### v2

# MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization

Abstract.
Prompt engineering is essential for fully leveraging large language models (LLMs), yet existing optimization methods often follow a single trajectory, leading to limited adaptability, gradient conflicts, and high computational overhead. We propose MAPGD (Multi-Agent Prompt Gradient Descent), a novel framework that reconceptualizes prompt optimization as a collaborative multi-agent process. MAPGD incorporates specialized agents focusing on orthogonal refinement dimensions—such as instruction clarity, example selection, format structuring, and stylistic adaptation—and coordinates their contributions through semantic gradient embedding, conflict detection, and fusion. To further enhance robustness and stability, MAPGD proposes two new mechanisms: Hypersphere-Constrained Gradient Clustering (HCGC), which enforces angular margin constraints to ensure intra-cluster compactness and inter-cluster separation, and Channel-Adaptive Agent Weighting (CAAW), which dynamically reweights agent contributions based on validation performance. Experiments on classification and reasoning benchmarks demonstrate that MAPGD consistently outperforms single-agent and random baselines in both accuracy and efficiency. Ablation studies validate the effectiveness of gradient fusion, agent specialization, and conflict resolution. Together, these contributions establish MAPGD as a unified, gradient-inspired, and interpretable framework for robust prompt optimization with theoretical convergence guarantees.
Key words and phrases:
Prompt Optimization, Multi-Agent Systems, Large Language Models, Gradient Descent
{CCSXML}
[…]
1. Introduction
Figure 1. Overview of the Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization.
[…]
Overview of the Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization.
Large Language Models (LLMs), trained on vast webscale corpora, have recently demonstrated remarkable generalization capabilities across diverse natural language processing (NLP) tasks Achiam et al. (2023); Bubeck et al. (2023). A central factor influencing their performance is the design of prompts, which serve as the primary interface for guiding model behavior. However, prompt construction remains largely a manual, trial-and-error process that demands substantial human labor Jiang et al. (2022) and domain expertise Reynolds and McDonell (2021); Zamfirescu-Pereira et al. (2023). This inefficiency highlights the urgent need for automatic or semi-automatic strategies to systematically generate or refine prompts. Such approaches can reduce manual burden, improve task performance, and provide interpretable optimization pathways.
Recent research has explored this problem through two primary directions. One line of work introduces auxiliary models or differentiable prompt embeddings to approximate gradient-based optimization Qin and Eisner (2021); Deng et al. (2022). Yet, these methods often assume access to the internal states or parameters of LLMs Shin et al. (2020); Lester et al. (2021), which is impractical in real-world API-only scenarios. Another line of work employs reinforcement learning or feedback-driven editing to discretely manipulate prompts Zhang et al. (2022); Zhou et al. (2022). While effective in some cases, these algorithms may produce incoherent prompts, incur high computational costs, or rely on unguided Monte Carlo search over the semantic space. Collectively, these limitations highlight the need for more robust, interpretable, and scalable frameworks for prompt optimization.
Addressing this gap, Pryzant et al. Pryzant et al. (2023) proposed ProTeGi, a nonparametric approach that mirrors gradient descent in textual space. ProTeGi generates natural language “gradients” from model errors and iteratively edits prompts in semantically opposite directions. To balance exploration and exploitation, it further integrates beam search and bandit-based candidate selection, achieving significant gains over earlier baselines. Nevertheless, ProTeGi operates under a single-agent paradigm, which restricts the diversity of optimization signals and leaves unresolved conflicts among competing refinement directions.
To overcome these limitations, we propose MAPGD (Multi-Agent Prompt Gradient Descent), a novel framework that reconceptualizes prompt optimization as a collaborative multi-agent process(Figure 1). MAPGD introduces a set of specialized agents, each dedicated to distinct refinement dimensions—such as instruction clarity, example selection, format structuring, and stylistic adaptation. To reconcile heterogeneous improvement signals, MAPGD employs a semantic gradient coordinator that embeds textual gradients into a shared vector space, enabling conflict detection, clustering, and fusion. In addition, we propose two novel mechanisms to further enhance robustness and stability: Hypersphere-Constrained Gradient Clustering (HCGC), which introduces angular margin constraints on the unit hypersphere to ensure compact intra-cluster coherence and inter-cluster separation, and Channel-Adaptive Agent Weighting (CAAW), which dynamically reweights agent contributions according to validation performance, amplifying effective optimization channels while suppressing noisy ones. Together, these innovations endow MAPGD with improved robustness, stability, and interpretability, enabling it to consistently outperform single-agent baselines such as ProTeGi across diverse optimization scenarios.
Beyond empirical improvements, MAPGD also provides a theoretical foundation for convergence. Under mild stochastic approximation assumptions, we prove that the multi-agent gradient coordination and fusion mechanisms preserve the convergence properties of classical gradient descent, yielding almost sure convergence to a local optimum at a sublinear rate of $\mathcal{O}(1/\sqrt{T})$, where $T$ denotes the number of iterations. This result establishes that, despite operating in a discrete, textual prompt space and coordinating multiple heterogeneous agents, MAPGD maintains stable and predictable optimization dynamics, bridging practical prompt refinement with rigorous theoretical guarantees.
Figure 2. Workflow of MAPGD. Starting from an initial prompt $p$, specialized agents propose diverse pseudo-gradients $\{g,g^{\prime},g^{\prime\prime},g^{\prime\prime\prime},\ldots\}$ (top). These gradients are semantically embedded and projected onto the unit hypersphere, then clustered with angular-margin constraints to form coherent groups (middle). Clustered directions are fused via LLM-driven synthesis into composite gradients $G$, which generate refined candidate prompts $\{P_{1},P_{2},\ldots,P_{n}\}$ (bottom). The best candidates are iteratively selected and fed back to continue the prompt optimization loop.
[…]
Workflow of MAPGD. Starting from an initial prompt $p$, specialized agents propose diverse pseudo-gradients $\{g,g^{\prime},g^{\prime\prime},g^{\prime\prime\prime}$…\} (top). These gradients are semantically embedded and projected onto the unit hypersphere, then clustered with angular-margin constraints to form coherent groups (middle). Clustered directions are fused via LLM-driven synthesis into composite gradients $G$, which generate refined candidate prompts $\{P_{1},P_{2},\ldots,P_{n}\}$ (bottom). The best candidates are iteratively selected and fed back to continue the prompt optimization loop.
2. Related Work
Prompt Learning and Optimization.
Early efforts relied on manual engineering, which lacks scalability. Automated methods include reinforcement learning Deng et al. (2022), evolutionary search Fernando et al. (2023), and Bayesian optimization Sahoo et al. (2025). Continuous prompt tuning (e.g., prefix tuning Li and Liang (2021), soft prompts Lester et al. (2021)) optimizes embeddings but sacrifices interpretability. MAPGD instead pursues interpretable natural language optimization with structured feedback.
Multi-Agent Collaboration.
Multi-agent systems are well-studied in RL, distributed AI, and game theory. In NLP, multi-agent debate Liang et al. (2023); Du et al. (2023), cooperative reasoning, and collaborative generation Hong et al. (2024) show that role specialization outperforms monolithic optimization. MAPGD applies this principle by assigning agents to distinct prompt dimensions and coordinating them via gradient fusion.
Gradient-Inspired Optimization.
ProTeGi Pryzant et al. (2023) and similar approaches approximate prompt gradients through self-feedback. While effective, single-agent descent struggles with limited signal diversity and inherent conflict resolution challenges Xiang et al. (2025). A monolithic agent can generate self-contradictory pseudo-gradients, as different refinement dimensions are not explicitly disentangled. For instance, to correct a misclassification, a single agent might propose a gradient to add more illustrative examples, while an implicit counter-gradient aims to make the instructions more concise. This creates a direct conflict in prompt length and structure, potentially leading to suboptimal or oscillating optimization trajectories. MAPGD extends this paradigm by embedding gradients semantically, explicitly detecting such conflicts, and fusing them coherently through LLM-driven synthesis.
Geometric and Adaptive Mechanisms.
Contrastive learning and angular margin constraints are effective in representation learning. Wang and Isola formalize how contrastive losses enforce alignment and uniformity, providing a theoretical lens Wang and Isola (2020). This principle of hyperspherical separation to enhance discriminability spans domains. In face recognition, margin-based softmax losses such as SphereFace Liu et al. (2017), CosFace Wang et al. (2018), and ArcFace Deng et al. (2019) enforce angular margins between identities. Chen et al. Chen et al. (2021) introduce a distance polarization regularizer for large-margin contrastive learning, pushing inter-cluster samples apart while keeping intra-cluster representations tight. Inspired by these advances, our Hypersphere-Constrained Gradient Clustering (HCGC) enforces intra-cluster compactness and inter-cluster separation for semantic gradients, complemented by Channel-Adaptive Agent Weighting (CAAW) adjusting agent contributions. Together, these modules enhance multi-agent gradient descent with geometry-aware clustering and adaptive coordination.
In summary, MAPGD unifies prompt learning, multi-agent collaboration, gradient descent, and geometry-inspired optimization into a scalable and interpretable framework.
3. Methodology
3.1. Framework Overview
MAPGD formulates prompt optimization as a hybrid discrete–continuous gradient descent process in the natural language space. Unlike continuous embedding–based methods such as prefix or soft prompt tuning Li and Liang (2021); Lester et al. (2021), MAPGD explicitly manipulates interpretable textual prompts while leveraging gradient-inspired feedback for iterative refinement (Figure 2).
At each iteration, specialized agents analyze misclassified or suboptimal examples and produce pseudo-gradients in natural language. These gradients are semantically embedded, clustered, and adaptively fused to mitigate conflicts and ensure coherent optimization. The resulting fused gradients are applied to generate successor prompts, which are then evaluated and filtered under limited budgets. Iterative repetition yields a dynamically evolving prompt trajectory that balances exploration of diverse directions with exploitation of promising refinements.
Formally, the optimization objective is defined as:
$F(p)=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\ell(M(x;p),y)\right],$ (1)
where $\ell(\cdot,\cdot)$ denotes a task-specific loss function. The goal is:
$p^{*}=\arg\min_{p}F(p).$ (2)
Unlike stochastic gradient descent (SGD), where gradients are continuous, MAPGD constructs textual pseudo-gradients:
$\nabla F(p^{(t)})\approx g^{(t)},$ (3)
which act as semantic analogues of numerical gradients.
3.2. Specialized Prompt Agents
Each agent specializes in a distinct dimension of prompt optimization, analogous to orthogonal gradient directions. For example: $A_{1}$: instruction clarity ($g_{1}$), $A_{2}$: example selection ($g_{2}$), $A_{3}$: format specification ($g_{3}$), $A_{4}$: stylistic refinement ($g_{4}$), and optionally a generic agent for broad improvements. At iteration $t$, the gradient set is:
$G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}.$
This decomposition supports parallel exploration of diverse refinements, alleviating local minima issues inherent in single-agent approaches.
3.3. Hypersphere-Constrained Gradient Clustering
Inspired by hyperspherical contrastive learning Wang et al. (2018); Liu et al. (2017), our HCGC mechanism is designed to enhance the semantic consistency of multi-agent gradients by enforcing intra-cluster compactness and inter-cluster separation in a normalized hypersphere space. This approach addresses the challenge of conflicting pseudo-gradients generated by heterogeneous agents, ensuring that semantically coherent optimization directions are emphasized while contradictory ones are disentangled. The overall process and the geometric intuition of our method are illustrated in Figure 3.
Gradient Embedding on Hypersphere.
Given a set of agent-generated gradients $G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}$, each gradient is encoded into a d-dimensional vector $v_{k}^{(t)}=\phi(g_{k}^{(t)})$, where $\phi$ denotes a pre-trained language encoder. For this purpose, we utilize a Sentence Transformer model, such as all-MiniLM-L6-v2, which provides compact and semantically informative embeddings. To eliminate scale variance, we normalize each embedding onto the unit hypersphere:
$\hat{v}_{k}^{(t)}=\frac{v_{k}^{(t)}}{\|v_{k}^{(t)}\|},\quad\hat{v}_{k}^{(t)}\in\mathbb{S}^{d-1}.$ (4)
The similarity between two gradients is then computed as the cosine similarity. Gradients are considered to be in conflict if their angular distance exceeds a predefined threshold, $\theta_{conflict}$:
$\mathrm{sim}(\hat{v}_{i},\hat{v}_{j})=\hat{v}_{i}^{\top}\hat{v}_{j}=\cos(\Delta(\hat{v}_{i},\hat{v}_{j})),$ (5)
where $\Delta(\cdot,\cdot)$ denotes the angular distance on the hypersphere.
Figure 4 illustrates a concrete example in the sentiment classification task, where different agents generate multiple pseudo-gradients. While some gradients form semantically coherent clusters (intra-cluster compactness), others introduce conflicts that must be separated through angular margin constraints.
Angular Margin Constraint.
To mitigate conflicts and ensure clear separation between semantic clusters, HCGC enforces an angular margin constraint on the unit hypersphere. Let $\hat{v}_{k}$ denote a normalized gradient vector, $u_{i}$ the centroid of its assigned cluster $i$, and $u_{j}$ the centroid of any other cluster $j$. We define the corresponding angles as
$\alpha=\Delta(\hat{v}_{k},u_{i}),\quad\beta=\Delta(\hat{v}_{k},u_{j}).$
A standard clustering assignment only requires $\alpha<\beta$. However, to enforce stronger separation, we introduce a margin scaling factor $n\geq 1$:
$n\cdot\alpha<\beta.$
Since $\cos(\theta)$ is monotonically decreasing for $\theta\in[0,\pi]$, this condition is equivalently expressed in terms of cosine similarity:
$\cos(n\cdot\alpha)>\cos(\beta),\quad n\geq 1.$ (6)
This constraint ensures that $\hat{v}_{k}$ is not only closer to its own centroid $u_{i}$, but also significantly more aligned with it than with any other centroid $u_{j}$. Geometrically, as illustrated in Figure 3 (c), this confines $\hat{v}_{k}$ to a scaled angular neighborhood around $u_{i}$, effectively creating a “cone of acceptance” narrower than the standard Voronoi region on the hypersphere. By enforcing this margin, HCGC simultaneously enhances inter-cluster separation and intra-cluster compactness, yielding more stable and reliable gradient fusion. During clustering, any gradient $\hat{v}_{k}$ violating this condition with respect to some $u_{j}$ is considered for reassignment to a cluster where the margin holds; otherwise, it remains in its original cluster.
Figure 3. Details of Hypersphere-Constrained Gradient Clustering
[…]
Details of Hypersphere-Constrained Gradient Clustering: Intra-cluster Compactness and Inter-cluster Repulsion, Original Boundary, Angular Margin
Optimization Objective.
Analogous to hyperspherical contrastive learning, the clustering objective combines intra-cluster compactness and inter-cluster repulsion:
$\mathcal{L}_{\text{HCGC}}=-\log\frac{\exp(\cos(\Delta(\hat{v}_{i},\hat{v}_{j}))/\tau)}{\sum_{k\in\mathcal{N}(i)}\exp(\cos(\Delta(\hat{v}_{i},\hat{v}_{k}))/\tau)},$ (7)
where $(\hat{v}_{i},\hat{v}_{j})$ is a positive pair from the same cluster, $\mathcal{N}(i)$ denotes negatives from other clusters, and $\tau$ is a temperature factor. By optimizing $\mathcal{L}_{\text{HCGC}}$, we obtain semantically well-aligned clusters of gradients, ensuring coherent downstream fusion.
After clustering, gradients within each semantically coherent group are fused into a single representative gradient. This fusion is typically performed by an LLM, guided by the original suggestions. To maintain diversity and avoid redundancy among the fused gradients, a semantic diversity filter is applied, discarding suggestions that are too similar to already selected ones, controlled by a diversity threshold $\theta_{diversity}$.
Figure 4. Illustration of agent-generated pseudo-gradients in a sentiment classification task. Blue gradients form a coherent cluster under the Instruction Specialist, while the red gradient represents a conflicting update. Green gradients from the Example Curator constitute another semantically compact cluster. This motivates the need for hypersphere-constrained clustering to enforce intra-cluster compactness and inter-cluster separation.
[…]
Illustration of agent-generated pseudo-gradients in a sentiment classification task. Blue gradients form a coherent cluster under the Instruction Specialist, while the red gradient represents a conflicting update. Green gradients from the Example Curator constitute another semantically compact cluster. This motivates the need for hypersphere-constrained clustering to enforce intra-cluster compactness and inter-cluster separation.
3.4. Channel-Adaptive Agent Weighting
While HCGC ensures structural consistency across gradients, gradients within the same cluster may still differ in reliability due to heterogeneous agent expertise. To address this, we propose Channel-Adaptive Agent Weighting (CAAW), which dynamically calibrates agent contributions according to their historical effectiveness, akin to channel-wise recalibration in neural representations Hu et al. (2018).
Adaptive Weight Assignment.
Let $s_{k}^{(t)}$ denote the validation performance gain associated with gradient $g_{k}^{(t)}$. The adaptive weight of each gradient is computed via a softmax normalization:
$w_{k}^{(t)}=\frac{\exp(\lambda s_{k}^{(t)})}{\sum_{j\in C}\exp(\lambda s_{j}^{(t)})},$ (8)
where $C$ is the current cluster, and $\lambda$ controls the sharpness of weighting. This formulation emphasizes historically effective agents while down-weighting less reliable ones.
Gradient Fusion.
The final fused gradient is expressed as:
$g_{\text{fused}}^{(t)}=\Psi\left(\sum_{k\in C}w_{k}^{(t)}g_{k}^{(t)}\right),$ (9)
where $\Psi(\cdot)$ denotes an LLM-driven synthesis function that integrates multiple textual directions into a coherent refinement.
Robustness and Stability.
By recalibrating agent contributions at the channel level, CAAW suppresses redundant or conflicting updates while amplifying stable and discriminative directions. Theoretically, this reduces variance in the optimization trajectory and prevents over-amplification of noisy gradients, thereby ensuring convergence to semantically robust prompt candidates.
3.5. Candidate Generation and Selection
Fused gradients generate successor prompts $\{p^{\prime}_{1},\dots,p^{\prime}_{n}\}$. MAPGD applies lightweight filtering and optional bandit-based evaluation Audibert and Bubeck (2010) to control computational overhead. However, unlike prior work that heavily relies on bandit selection, our framework emphasizes geometry-aware clustering (HCGC) and adaptive agent weighting (CAAW) as the primary mechanisms ensuring robustness and interpretability.
3.6. Theoretical Convergence (Informal)
We provide a convergence analysis of MAPGD in Appendix C. Here we summarize the key assumptions: (i) bounded variance of agent pseudo-gradients, (ii) Lipschitz continuity of the loss function, and (iii) unbiased sampling from bandit-based candidate selection. Under these conditions, MAPGD inherits the convergence properties of stochastic gradient descent and achieves almost sure convergence to a local optimum at a sublinear rate of $\mathcal{O}(1/\sqrt{T})$. Intuitively, HCGC reduces variance by enforcing geometric separation, CAAW emphasizes historically reliable agents, and bandit selection maintains unbiased exploration, together ensuring stable optimization dynamics.
4. Experiments
We design comprehensive experiments to evaluate the effectiveness, efficiency, and robustness of the proposed MAPGD framework. To ensure comparability, we reproduce the experimental setup of ProTeGi under identical conditions and extend the evaluation with ablation studies and token consumption analysis. In addition, we include a real-world case study on prompt optimization for text generation (see Appendix D), which highlights the applicability of MAPGD beyond controlled benchmarks.
4.1. Datasets
To ensure comparability with ProTeGi, we evaluate MAPGD on the same four benchmark NLP classification tasks that span diverse domains and languages:
  • Jailbreak Shen et al. (2024): a novel task to determine whether a user input to an LLM continuation API constitutes a jailbreak attempt. Jailbreak attacks are defined as user strategies that aim to make the model violate its own safety policies, such as generating harmful content or exposing hidden instructions. The dataset contains 1306 multilingual examples with human-annotated jailbreak labels.
  • Ethos Mollas et al. (2022): an English hate speech detection dataset with 997 online comments annotated as hateful or non-hateful.
  • LIAR Wang (2021): a large-scale dataset for fake news detection, consisting of 4,000 short statements labeled with ground-truth veracity, along with context information.
  • Sarcasm Farha and Magdy (2020): an Arabic sarcasm detection dataset with 10,000 online comments, labeled for the presence or absence of sarcasm.
Beyond classification, we further evaluate on three arithmetic reasoning datasets that require multi-step logical inference and symbolic manipulation:
  • GSM8k Cobbe et al. (2021): a widely used benchmark of 8,500 grade-school math problems that require step-by-step reasoning.
  • AQUARAT Ling et al. (2017): a dataset of algebraic word problems designed to test symbolic reasoning and program induction.
  • SVAMP Patel et al. (2021): a benchmark focusing on simple arithmetic word problems with diverse linguistic variations, serving as a testbed for robustness to paraphrasing.
In addition, we include a real-world text generation case study in the Appendix to highlight MAPGD’s applicability beyond controlled benchmarks.
4.2. Experimental Setup
For each dataset, we randomly sample 50 examples for development and 150 for testing. All reported results are averaged over three independent runs to mitigate variance. The evaluation metric is binary F1 score on the test set, computed by max-pooling over the final beam of candidate prompts.
Unless otherwise stated, all experiments are conducted with the January 2023 version of gpt-3.5-turbo through the Azure OpenAI API. The temperature is set to 0.0 for classification tasks to ensure deterministic predictions.
To ensure a fair comparison, all methods—including baselines and MAPGD—use the same initial prompts, training data, and random seeds. The iteration budget is fixed at $T=10$ across all experiments. We adopt default parameter values without additional hyperparameter search, to ensure that performance differences are attributable to algorithmic design rather than parameter tuning.
For our MAPGD implementation, we instantiate the framework with specific hyperparameters introduced in Section 3.3. We use the all-MiniLM-L6-v2 Sentence Transformer as the gradient encoder $\phi$. The conflict threshold $\theta_{conflict}$ is set to 0.3, and we use a cluster similarity threshold of 0.7 for grouping gradients. The diversity threshold $\theta_{diversity}$ is set to 0.7 to ensure varied candidate prompts. Bandit-based selection adopts a UCB strategy with a budget of 80 evaluations.We set the CAAW weighting parameter $\lambda$ to 1 (see Appendix A for a sensitivity analysis on different values).
Accordingly, all experiments reported in Section 4.4 were conducted with these modules enabled, unless explicitly ablated.
4.3. Baselines
To evaluate the effectiveness of MAPGD, we compare it against a set of non-parametric prompt optimization methods, following the setup of ProTeGi Pryzant et al. (2023), and additionally include ProTeGi itself as a strong baseline. Specifically, we consider:
  • ProTeGi: the original prompt gradient descent framework, where a single agent iteratively generates pseudo-gradients and candidate prompts, with bandit-based selection. This serves as our primary baseline for comparison.
  • Monte-Carlo (MC) Zhou et al. (2022): an iterative but directionless Monte Carlo search over the prompt space. For fairness, we match the number of samples per candidate to the successors generated by MAPGD.
  • Reinforcement Learning (RL) Recent approaches such as GrIPS Prasad et al. (2022) and TEMPERA Zhang et al. (2022) formulate prompt optimization as a reinforcement learning problem. In these methods, the prompt text is first segmented into phrases, and the search space is explored via phrase-level edit operations, including addition, paraphrasing, swapping, and deletion.
  • AutoGPT: an open-source autonomous agent system that improves prompts through self-directed feedback loops. We configure AutoGPT with the same number of examples and errors as MAPGD, running for the same number of optimization steps.
For arithmetic reasoning, we additionally compare against three competitive baselines, all reported with GPT-3.5-Turbo in a zero-shot setting:
  • InsZero Chen et al. (2023): a zero-shot prompting method with instruction tuning.
  • Instinct Lin et al. (2024): an instruction-following baseline that improves prompting robustness.
  • PromptWizard Agarwal et al. (2024): a state-of-the-art prompt synthesis framework that iteratively enriches prompts with intermediate reasoning steps and examples.
This collection of baselines allows us to evaluate MAPGD not only against single-agent gradient-based optimization, but also against strong arithmetic reasoning methods optimized for chain-of-thought style prompting.
4.4. Experimental Results
Figure 5 summarizes the overall performance across the four benchmark classification datasets. Our results demonstrate that MAPGD consistently achieves the best performance under varying evaluation budgets, outperforming all baselines by a notable margin.
On the Jailbreak dataset, MAPGD attains an average F1 of 0.86, surpassing ProTeGi (0.82) and MC (0.76), while substantially improving over RL (0.67) and AutoGPT (0.62). The improvement margin remains stable across different budgets, reaching up to +5.5% over ProTeGi and +10.9% over MC at 30 evaluations.
For Ethos, MAPGD pushes performance from the strong baseline of 0.93 ($p_{0}$) to 0.98, showing a consistent +1.5% advantage over ProTeGi and +4.0% over MC. These improvements highlight MAPGD’s ability to leverage gradient fusion even in relatively saturated tasks.
On the more challenging LIAR dataset, MAPGD achieves 0.71, representing the largest relative gains: +8.0% over ProTeGi, +11.0% over MC, and a substantial +17.0% over the initial prompt. This suggests that multi-agent collaboration and semantic clustering are particularly beneficial in noisy, fact-checking tasks.
For Sarcasm, MAPGD steadily improves from 0.84 ($p_{0}$) to 0.91, outperforming ProTeGi (0.87) and MC (0.85). Interestingly, AutoGPT sometimes reduces performance below the initial prompt, echoing earlier findings that unguided agent feedback may destabilize optimization.
Beyond classification, we further evaluate MAPGD on arithmetic reasoning benchmarks, which require detailed multi-step inference. Table 1 compares MAPGD against three strong baselines—InsZero, Instinct, and PromptWizard—on GSM8k, AQUARAT, and SVAMP. MAPGD achieves consistent improvements across all datasets, with particularly notable gains on GSM8k (93.5%), surpassing the previous best by +3.5 points. On AQUARAT and SVAMP, MAPGD reaches 60.3% and 84.1%, respectively, representing stable improvements over the strongest baselines. These results confirm that MAPGD not only excels in classification but also transfers effectively to challenging reasoning tasks.
Overall, MAPGD improves over ProTeGi and MC by 3–7% on average for classification, while maintaining strong gains on arithmetic reasoning datasets. These results validate our hypothesis that multi-agent specialization and hypersphere-constrained clustering provide more robust and adaptive optimization compared to single-agent or rule-based approaches.
Table 1. Arithmetic reasoning accuracy (%) on GSM8k, AQUARAT, and SVAMP. Best results in bold.
Method GSM8k AQUARAT SVAMP
InsZero 74.2 54.3 79.5
Instinct 74.5 54.7 81.0
PromptWizard 90.0 58.2 82.3
MAPGD 93.5 60.3 84.1
Figure 5. Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks.
[…]
Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks.
4.5. Ablation Study
To precisely attribute the performance gains of MAPGD, we perform a comprehensive ablation study (see Table 2 for the detailed configurations and results). Beyond standard ablation variants (e.g., No Gradient Fusion, Single-Agent Only, No Specialization), we design targeted experiments specifically to isolate the contributions of the two proposed modules HCGC and CAAW, as well as to evaluate resource efficiency and robustness.
Ablation configurations.
We evaluate the following variants:
  1. (1)
    MAPGD (full): our full method with HCGC and CAAW.
  2. (2)
    No Gradient Fusion: disable the fusion stage; agents act independently and their suggestions are applied without semantic clustering.
  3. (3)
    Single-Agent Only: only one specialized agent (instruction specialist) is used.
  4. (4)
    No Specialization: agent roles are randomized each iteration (no fixed expertise).
  5. (5)
    w/o HCGC: keep clustering but remove the hypersphere angular-margin constraint. This reduces to standard K-means clustering with cosine similarity, which acts as an off-the-shelf baseline for comparison.
  6. (6)
    w/o CAAW: replace Channel-Adaptive Agent Weighting with simple uniform averaging of agent gradients.
  7. (7)
    w/o HCGC+CAAW: use standard K-means clustering with cosine similarity (no angular margin) and replace adaptive weighting with uniform fusion.
Table 2. Ablation study: classification performance (F1 score, mean $\pm$ std over 3 runs). Best results in bold.
Method Jailbreak Ethos LIAR Sarcasm
MAPGD (full) 0.88 $\pm$ 0.01 0.98 $\pm$ 0.02 0.71 $\pm$ 0.01 0.91 $\pm$ 0.01
No Gradient Fusion 0.83 $\pm$ 0.02 0.92 $\pm$ 0.02 0.67 $\pm$ 0.02 0.86 $\pm$ 0.02
Single Agent Only 0.80 $\pm$ 0.03 0.87 $\pm$ 0.03 0.62 $\pm$ 0.02 0.82 $\pm$ 0.03
No Specialization 0.80 $\pm$ 0.02 0.90 $\pm$ 0.02 0.65 $\pm$ 0.02 0.84 $\pm$ 0.02
w/o HCGC 0.84 $\pm$ 0.02 0.94 $\pm$ 0.02 0.68 $\pm$ 0.02 0.87 $\pm$ 0.02
w/o CAAW 0.85 $\pm$ 0.02 0.95 $\pm$ 0.02 0.69 $\pm$ 0.02 0.88 $\pm$ 0.02
w/o HCGC + CAAW 0.82 $\pm$ 0.03 0.91 $\pm$ 0.03 0.66 $\pm$ 0.02 0.83 $\pm$ 0.02
4.6. Token Consumption Analysis
To further assess the efficiency of MAPGD, we conducted a dedicated token consumption experiment, following the same setup as in Section 4.2. In this experiment, we recorded the number of tokens used in every API call, where the token cost of a single call is defined as the sum of input tokens and output tokens. For each method, we accumulated token usage across all iterations and runs, and report the total tokens, average tokens per iteration, and total number of LLM calls. In addition, we report performance efficiency as the ratio of average F1 score to total token usage.
Table 3 presents the results, comparing MAPGD with the ProTeGi. MAPGD not only achieves higher F1 scores, but also requires fewer tokens and API calls overall. Specifically, MAPGD reduces the average token cost per iteration from roughly 25.6k to 23.6k, corresponding to a $\sim$8% saving in token usage. At the same time, MAPGD improves the performance-per-token efficiency by more than 10%. These findings indicate that multi-agent collaboration with HCGC and CAAW enables more effective use of query budgets, reducing redundancy while preserving accuracy.
Table 3. Token consumption comparison between ProTeGi and MAPGD (averaged over 3 runs). MAPGD achieves lower token usage and higher efficiency.
Metric ProTeGi MAPGD (full)
Total Tokens 256k 236k
Calls 962 643
Avg F1 0.83 0.87
F1 / Token ($\times 10^{-6}$) 3.24 3.69
4.7. Agent Number Sensitivity
To better justify the design choice of using four specialized agents, we conduct a sensitivity analysis by varying the number of agents $N\in\{2,4,6\}$. For $N=2$, we retain only the Instruction Specialist and Example Curator; for $N=6$, we extend the framework with two additional stylistic and format-focused agents. All other settings follow Section 4.2.
Figure 6 shows the average F1 score across the four datasets. We observe that performance improves substantially when moving from $N=2$ to $N=4$, confirming the benefit of multi-agent specialization. Increasing to $N=6$ yields only marginal gains while incurring higher token cost, suggesting diminishing returns. These results indicate that four agents provide a favorable balance between effectiveness and efficiency, validating our default configuration.
Overall, these findings suggest that an excessively small agent set cannot comprehensively capture different prompt dimensions, whereas an overly large set introduces redundancy and potential conflicts. A moderate number of agents, four in our case, provides the best trade-off.
Figure 6. Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
[…]
Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
4.8. Experimental Insights
Our experimental evaluation provides several key insights:
  • Effectiveness of multi-agent collaboration. Across all four benchmark tasks, MAPGD consistently outperforms single-agent ProTeGi and other baselines. The largest gains are observed on LIAR and Jailbreak, where conflicting gradient signals are common. This demonstrates that multi-agent specialization, combined with semantic clustering, can effectively resolve gradient conflicts and guide prompt optimization toward more robust improvements.
  • Contributions of HCGC and CAAW. The ablation study confirms that both proposed modules are indispensable. Removing HCGC reduces the ability to enforce semantic compactness and separation, leading to less stable optimization. Disabling CAAW degrades robustness by failing to downweight weaker agents. When both modules are removed, MAPGD collapses close to single-agent performance, highlighting their complementary roles.
  • Efficiency in token usage. Despite employing multiple agents, MAPGD achieves lower overall token consumption and fewer API calls than ProTeGi. This efficiency stems from clustering redundant gradients and adaptively weighting agents, which reduce unnecessary evaluations. As a result, MAPGD achieves over 10% higher performance-per-token efficiency, indicating that it not only improves accuracy but also makes more economical use of query budgets.
  • Robustness across tasks. The improvements are consistent across datasets of varying size, domain, and language, suggesting that MAPGD generalizes well to heterogeneous NLP tasks. Furthermore, the case study in the Appendix illustrates its applicability to open-ended generation, extending beyond controlled classification benchmarks.
  • Transferability to reasoning tasks. MAPGD also demonstrates strong performance on arithmetic reasoning datasets, surpassing competitive baselines such as PromptWizard. Its consistent improvements on GSM8k, AQUARAT, and SVAMP highlight that the proposed gradient-inspired multi-agent optimization is not limited to classification, but can effectively enhance chain-of-thought style reasoning as well.
In summary, these insights show that MAPGD achieves a favorable balance between performance and efficiency, while its modular design (HCGC + CAAW) provides both effectiveness and robustness in diverse prompt optimization scenarios.
5. Discussion
MAPGD advances prompt optimization by combining gradient-inspired reasoning with multi-agent collaboration. Its two key mechanisms—Hypersphere-Constrained Gradient Clustering and Channel-Adaptive Agent Weighting enable more structured, interpretable, and efficient optimization.
Modularity. By separating gradient generation, clustering, weighting, and candidate selection, MAPGD clarifies which components drive performance, supporting transparent analysis and principled extensions.
Conflict resolution. HCGC enforces angular margin constraints on the hypersphere, promoting intra-cluster compactness and inter-cluster separation. This directly addresses conflicts among agent-generated gradients, common in tasks such as fake news detection or jailbreak classification, and parallels gradient surgery in multi-task learning but for discrete prompts.
Adaptive collaboration. CAAW allocates influence based on agent reliability, reducing noise from weaker gradients while amplifying consistent signals. This is useful when tasks emphasize different reasoning dimensions (e.g., semantic consistency vs. contextual grounding).
Budget-awareness. With lightweight bandit-based selection, MAPGD improves performance-per-token efficiency, achieving higher accuracy with fewer redundant API calls.
Limitations. MAPGD depends on embedding quality and agent reliability; under strong domain shifts, clustering may be unstable and weighting may amplify spurious signals. Future work may strengthen semantic representations and improve agent calibration.
Future directions. Promising avenues include:
  1. (1)
    Cross-task generalization: training reusable agents with transferable clustering and weighting behaviors.
  2. (2)
    Human-in-the-loop fusion: leveraging expert oversight to guide clustering and validate weighting.
  3. (3)
    Hybrid optimization: integrating MAPGD with differentiable prompt tuning to balance interpretability and efficiency.
6. Conclusion
We presented MAPGD: Multi-Agent Prompt Gradient Descent, a novel framework for optimizing prompts in large language models. The key innovations of MAPGD are HCGC, which resolves conflicts through hypersphere-constrained clustering, and CAAW, which adaptively balances the contributions of specialized agents. Together, these mechanisms enable diverse agents to collaborate effectively while maintaining semantic alignment.
Empirically, MAPGD achieves consistent improvements over strong baselines such as ProTeGi, Monte Carlo search, reinforcement learning–based methods, and recent arithmetic reasoning frameworks. On four benchmark classification tasks, MAPGD improves F1 scores by 3–7% on average while reducing token consumption by $\sim$8%, leading to better performance-per-token efficiency. Beyond classification, MAPGD also demonstrates strong transferability to arithmetic reasoning datasets such as GSM8k, AQUARAT, and SVAMP, where it surpasses competitive baselines including PromptWizard. These results confirm that gradient-inspired multi-agent collaboration is broadly applicable across both discriminative and reasoning tasks.
Theoretically, we show that MAPGD achieves a sublinear convergence rate of $O(1/\sqrt{T})$ under standard stochastic approximation assumptions (see Appendix C for details). This result indicates that even in a discrete prompt space, the semantic gradient mechanism—stabilized by HCGC and CAAW—retains the efficiency of classical stochastic gradient descent.
While challenges remain, particularly regarding embedding robustness and LLM feedback reliability, MAPGD provides both a practical toolkit for efficient prompt optimization and a theoretical foundation linking discrete semantic updates with continuous optimization theory. Future work may extend HCGC and CAAW to multimodal prompts, incorporate human preference alignment, and develop cross-domain reusable agents.
In summary, MAPGD demonstrates that multi-agent collaboration, when combined with principled clustering and adaptive weighting, leads to more robust, interpretable, and resource-efficient prompt optimization, with demonstrated effectiveness across classification, reasoning, and beyond.
Appendix A CAAW Parameter Sensitivity
The Channel-Adaptive Agent Weighting module uses a temperature parameter $\lambda$ to control the emphasis on historically effective agents during gradient fusion. To evaluate the robustness of MAPGD with respect to $\lambda$, we conducted experiments on two representative datasets: LIAR(binary) and GSM8K (mathematical reasoning).
We tested three values of $\lambda$: 0.5, 1, and 2. Table 4 summarizes the test accuracy for each setting. The results indicate that moderate variations of $\lambda$ do not significantly affect the overall performance, supporting the choice of $\lambda=1$ in the main experiments.
Table 4. Test accuracy of MAPGD under different CAAW weighting parameter $\lambda$. Performance differences are minor, demonstrating robustness to $\lambda$.
$\lambda$ LIAR Accuracy GSM8K Accuracy
0.5 0.709 0.927
1 0.717 0.935
2 0.711 0.933
As shown in Table 4, the test accuracy exhibits only minor variations across the three $\lambda$ values, with a maximum difference of 0.008 on LIAR and 0.008 on GSM8K. These variations are small relative to the overall performance gains achieved by multi-agent gradient fusion, indicating that while $\lambda$ influences the sharpness of agent weighting, MAPGD’s performance remains stable for reasonable choices around the default setting of $\lambda=1$.
Appendix B Complexity and Parallelism
The computational bottleneck of MAPGD with HCGC and CAAW (see Algorithm 1) lies in LLM calls for gradient generation, cluster fusion, and prompt expansion, whereas embedding, clustering, similarity checks, and adaptive weighting involve comparatively lightweight vector operations. Multi-agent parallelization significantly reduces wall time, making the framework scalable for practical deployment. In the following, we analyze the per-iteration computational and memory costs, providing a detailed breakdown of the dominant operations and their respective complexities.
Notation.
[…]
Stage costs (time).
  1. (1)
    Agent gradient generation: Each of the $N$ agents generates $m$ gradients using LLMs, giving $O(N)$ LLM calls. Parallelization reduces wall time to $\approx\max t_{\text{LLM}}$.
  2. (2)
    HCGC embedding + conflict detection: Embedding each gradient costs $O(Gd)$; computing pairwise angular conflicts costs $O(G^{2}d)$. For small $G$ (e.g., $G\sim 16$), this is negligible compared with LLM calls.
  3. (3)
    HCGC clustering: KMeans clustering over $G$ gradients into $K$ clusters with $I$ iterations costs $O(GKId)$, again minor due to small $G$.
  4. (4)
    HCGC fusion (LLM): Clusters with multiple gradients require one LLM call per cluster to fuse into a coherent gradient. Total LLM calls up to $O(K_{\text{merge}})$, parallelizable across clusters.
  5. (5)
    CAAW weighting and fusion: Per-cluster adaptive weighting is $O(G)$ vector operations, negligible compared to LLM fusion.
  6. (6)
    Prompt expansion + MC paraphrasing: Applying each fused $O(|\tilde{\mathcal{G}}|) gradient$ applications. MC paraphrasing for each variant: $O(|\tilde{\mathcal{G}}|sn_{\text{mc}})$ LLM calls.
  7. (7)
    Diversity filtering: Computing embeddings: $O(|\mathcal{C}|d)$; naive pairwise similarities: $O(|\mathcal{C}|^{2})$. For tens of candidate prompts, this remains negligible.
  8. (8)
    Bandit evaluation: Probing $K_{\text{eval}}$ arms over mini-batches of size $b$ for $T_{b}$ rounds: $O(K_{\text{eval}}bT_{b})$, significantly cheaper than exhaustive evaluation $O(|\mathcal{C}|\,|D_{\text{dev}}|)$.
Space complexity.
Text storage: $O(|\mathcal{C}|L_{\text{avg}})$. Embeddings: $O((G+|\mathcal{C}|)d)$, typically a few MB. Optional caches (unique prompts or intermediate LLM outputs) scale linearly with the number of prompts and clusters.
Algorithm 1 MAPGD with HCGC and CAAW
1:Initial prompt $p_{0}$, train data $D_{\text{train}}$, dev data $D_{\text{dev}}$, agents $\{A_{i}\}_{i=1}^{N}$, iterations $R$, beam width $k$
2:Initialize best prompt $p_{\star}^{(0)}\leftarrow p_{0}$, beam $B_{0}\leftarrow\{p_{0}\}$; initialize each agent prompt $A_{i}.p\leftarrow p_{0}$
3:for $t=1$ to $R$ do
4:  $M_{t}\leftarrow\textsc{SampleMiniBatch}(D_{\text{train}},b)$
5:  $\mathcal{G}_{t}\leftarrow\textsc{GenerateAgentGradients}(\{A_{i}\},M_{t})$ $\triangleright$ Alg. 2
6:  $\tilde{\mathcal{G}}_{t}\leftarrow\textsc{HCGC}(\mathcal{G}_{t})$ $\triangleright$ Alg. 3
7:  $F_{t}\leftarrow\textsc{CAAWFuse}(\tilde{\mathcal{G}}_{t},D_{\text{dev}})$ $\triangleright$ Alg. 4
8:  $\mathcal{C}_{t}\leftarrow\textsc{ExpandPrompts}(p_{\star}^{(t-1)},F_{t})$
9:  $B_{t},p_{\star}^{(t)}\leftarrow\textsc{BanditSelect}(\mathcal{C}_{t},D_{\text{dev}},k)$
[…]
12:  end if
13:end for
14:return Best prompt over $\{p_{\star}^{(1)},\dots,p_{\star}^{(t)}\}$
Algorithm 2 Multi-Agent Textual Gradient Generation
1:Agents $\{A_{i}\}$, mini-batch $M_{t}$, task $T$, predictor $\Pi$, per-agent error cap $e$, feedback count $m$
2:for all agent $A_{i}$ in parallel do
3:  $(\hat{y},y)$ pairs $\leftarrow T.\textsc{InferAndLabel}(A_{i}.p,M_{t},\Pi)$
4:  $E_{i}\leftarrow\textsc{SelectErrors}(\hat{y},y,e)$
5:  if $|E_{i}|=0$ then $E_{i}\leftarrow\textsc{DiverseSamples}(M_{t},e)$
6:  end if
7:  $raw_{i}\leftarrow\textsc{LLMGradientPrompt}(A_{i}.role,A_{i}.p,E_{i},m)$
8:  $g_{i}\leftarrow\textsc{ParseGradientBlocks}(raw_{i})$ $\triangleright$ Split by delimiters
9:end for
10:return $\mathcal{G}_{t}=\{(A_{i}.role,g_{i})\}_{i=1}^{N}$
Algorithm 3 Hypersphere-Constrained Gradient Clustering
1:Agent gradients $\mathcal{G}_{t}$
2:Embed and normalize: $\hat{v}_{k}=\phi(g_{k})/\|\phi(g_{k})\|$
3:Detect conflicts: $\mathcal{C}_{\text{conf}}\leftarrow\{(i,j):\cos^{-1}(\hat{v}_{i}^{\top}\hat{v}_{j})>\theta\}$
4:Cluster gradients: $\{S_{k}\}_{k=1}^{K}\leftarrow\textsc{KMeans}(\{\hat{v}_{k}\},K_{\max})$
5:for $k=1$ to $K$ do
6:  Apply angular margin: reassign gradients violating $\cos(n\cdot\Delta(\hat{v}_{i},c_{k}))>\cos(\Delta(\hat{v}_{i},c_{j}))$
7:  Fuse cluster $S_{k}$ via LLM: $f_{k}\leftarrow\textsc{LLMFuse}(S_{k},\mathcal{C}_{\text{conf}})$
8:end for
9:return Fused clusters: $\tilde{\mathcal{G}}_{t}=\{f_{1},\dots,f_{K}\}$
Algorithm 4 Channel-Adaptive Agent Weighting and Fusion
1:Clustered gradients $\tilde{\mathcal{G}}_{t}$, dev data $D_{\text{dev}}$, temperature $\lambda$
2:for all cluster $\tilde{\mathcal{G}}_{t}$ do
3:  Compute per-gradient validation gain $s_{i}$ on $D_{\text{dev}}$
4:  Assign adaptive weight: $w_{i}=\frac{\exp(\lambda s_{i})}{\sum_{j\in S_{k}}\exp(\lambda s_{j})}$
5:  Fuse cluster gradients: $f_{k}=\Psi\Big(\sum_{i\in S_{k}}w_{i}g_{i}\Big)$
6:end for
7:return $F_{t}=\{f_{1},\dots,f_{K}\}$
Algorithm 5 SynchronizeAgents
1:Agents $\{A_{i}\}_{i=1}^{N}$, current best prompt $p_{\star}^{(t)}$
2:for each agent $A_{i}$ do
3:  $A_{i}.p\leftarrow p_{\star}^{(t)}$ $\triangleright$ Set the agent’s current prompt to the global best
4:  ResetGradientHistory($A_{i}$) $\triangleright$ Optional: clear outdated gradient history
[…]
6:end for
7:return Updated agents $\{A_{i}\}$
Appendix C Theoretical Analysis
In this section, we provide convergence guarantees for MAPGD under mild assumptions, following the stochastic approximation framework. Our goal is to bridge the gap between the continuous optimization theory of stochastic gradient descent (SGD) and the discrete prompt optimization carried out in MAPGD. We show that, despite operating in a structured and discrete search space, MAPGD achieves the same sublinear convergence rate of $O(1/\sqrt{T})$ in both convex and non-convex settings.
C.1. Assumptions
We begin with a set of assumptions standard in stochastic optimization but reinterpreted in the context of multi-agent prompt optimization.
  • (A1) Alignment (Unbiasedness). For some $\mu>0$, the stochastic semantic gradient $g^{(t)}$ maintains alignment with the true gradient:
    $\mathbb{E}\!\left[\langle g^{(t)},\nabla F(p^{(t)})\rangle\,\middle|\,p^{(t)}\right]\;\geq\;\mu\|\nabla F(p^{(t)})\|^{2}.$
    This reflects the role of semantic fusion: multi-agent aggregation reduces the chance of adversarial or noisy updates, ensuring progress along descent directions.
  • (A2) Bounded Second Moment. For constants $\rho,\sigma^{2}\geq 0$,
    $\mathbb{E}\!\left[\|g^{(t)}\|^{2}\,\middle|\,p^{(t)}\right]\;\leq\;\rho\|\nabla F(p^{(t)})\|^{2}+\sigma^{2}.$
    This captures the variance-control effect of the bandit-based selection mechanism, which prevents uncontrolled explosion of gradient magnitude.
  • (A3) Smoothness or Lipschitzness. For convex tasks, $F$ is $G$-Lipschitz with domain diameter $D$. For non-convex tasks, $F$ is $L$-smooth: $\|\nabla F(u)-\nabla F(v)\|\leq L\|u-v\|$.
C.2. Supporting Lemmas
We restate two standard lemmas, adapted to the MAPGD setting.
Lemma 1 (Convex Projection Inequality).
For convex $F$ with feasible set $\mathcal{P}$, the projected subgradient update
$p^{(t+1)}=\Pi_{\mathcal{P}}\!\left(p^{(t)}-\eta g^{(t)}\right)$
satisfies
$\|p^{(t+1)}-p^{*}\|^{2}\;\leq\;\|p^{(t)}-p^{*}\|^{2}-2\eta\langle g^{(t)},p^{(t)}-p^{*}\rangle+\eta^{2}\|g^{(t)}\|^{2}.$
Lemma 2 (Non-Convex Descent Lemma).
If $F$ is $L$-smooth, then for update $p^{(t+1)}=p^{(t)}-\eta g^{(t)}$, we have
$F(p^{(t+1)})\;\leq\;F(p^{(t)})-\eta\langle\nabla F(p^{(t)}),g^{(t)}\rangle+\frac{L}{2}\eta^{2}\|g^{(t)}\|^{2}.$
C.3. Main Results
Convex Convergence.
Suppose $F$ is convex, $G$-Lipschitz, and $\mathcal{P}$ has diameter $D$. Let $\bar{p}_{T}=\tfrac{1}{T}\sum_{t=1}^{T}p^{(t)}$. Under (A1)–(A2) and step size $\eta=\tfrac{D}{G\sqrt{T}}$, we obtain:
$\mathbb{E}[F(\bar{p}_{T})]-F(p^{*})\;=\;O\!\left(\frac{1}{\sqrt{T}}\right).$
Proof sketch. By Lemma 1 and convexity:
$F(p^{(t)})-F(p^{*})\leq\langle g^{(t)},p^{(t)}-p^{*}\rangle.$
Summing over $t=1,\ldots,T$ and applying (A1)–(A2), we bound the regret:
$\sum_{t=1}^{T}\mathbb{E}[F(p^{(t)})-F(p^{*})]\;\leq\;\frac{D^{2}}{2\eta}+\frac{\eta G^{2}T}{2}.$
Using Jensen’s inequality for $\bar{p}_{T}$ and optimizing $\eta$, we conclude the $O(1/\sqrt{T})$ rate.
Non-Convex Convergence.
Suppose $F$ is $L$-smooth and (A1)–(A2) hold. With constant step size $\eta=\Theta(1/\sqrt{T})$, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left[\|\nabla F(p^{(t)})\|^{2}\right]=O\!\left(\frac{1}{\sqrt{T}}\right).$
Proof sketch. Applying Lemma 2 and taking conditional expectation:
$\mathbb{E}[F(p^{(t+1)})]$ $\leq\mathbb{E}[F(p^{(t)})]-\eta\mu\mathbb{E}[\|\nabla F(p^{(t)})\|^{2}]$
$\quad+\tfrac{L}{2}\eta^{2}\left(\rho\mathbb{E}[\|\nabla F(p^{(t)})\|^{2}]+\sigma^{2}\right).$ (10)
Summing over $t=1\ldots T$ gives
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left[\|\nabla F(p^{(t)})\|^{2}\right]\;\leq\;\frac{2(F(p^{(1)})-F_{\inf})}{\mu T\eta}+\frac{L\sigma^{2}}{\mu}\eta.$
Balancing terms with $\eta=\Theta(1/\sqrt{T})$ yields the claimed rate.
C.4. Summary of Theoretical Guarantees
In this work, we bridge the gap between classical stochastic optimization, which assumes continuous parameter spaces, and the inherently discrete, structured nature of prompt optimization. Our convergence analysis for MAPGD does not require the discrete prompt space to be continuous. Instead, we show that the framework’s core mechanisms ensure that the optimization process satisfies the key conditions of stochastic approximation theory.
The argument unfolds along three main dimensions:
  • Directional Alignment (A1): Although textual pseudo-gradients are not true mathematical gradients, Hypersphere-Constrained Gradient Clustering aggregates semantically coherent suggestions from multiple agents into a fused gradient that statistically aligns with the true descent direction.
  • Variance Control (A2): Discrete prompt edits are prone to high variance, but the bandit-based candidate selection mechanism filters out unreliable or detrimental candidates. This effectively bounds the second moment of stochastic updates.
  • Smoothness Justification (A3): We assume the empirical loss function is Lipschitz or smooth in the semantic embedding space, even though the prompts themselves are discrete.
Given that these conditions are met, MAPGD inherits the convergence guarantees of classical stochastic gradient descent. Consequently, despite operating in a discrete textual space, MAPGD achieves a robust sublinear convergence rate of $O(1/\sqrt{T})$, providing a solid theoretical foundation for the observed empirical effectiveness of our multi-agent, geometry-aware approach to prompt optimization.
Appendix D Case Study: System Prompt Optimization
To further illustrate the applicability of MAPGD, we present a case study where our method is applied to optimize the system prompt of a large language model assistant. The original prompt is designed to support multi-source analysis and financial data interpretation for decision-making tasks. Using MAPGD, we refine the prompt to enhance robustness, accuracy, and interpretability by embedding explicit verification protocols, structured analysis guidelines, and risk prioritization frameworks. This example highlights how MAPGD can be deployed in practical LLM applications beyond benchmark datasets, particularly in domains where data authenticity, reliability, and interpretability are critical.
Original System Prompt. You are an AI assistant designed to process, analyze, and synthesize information from multiple sources in order to answer user questions, generate insights, and prepare detailed reports. You have specialized capabilities in financial data interpretation, knowledge retrieval, and multi-source analysis. You support both operational and strategic decision-making for Golden Section’s portfolio companies.
Optimized System Prompt via MAPGD. You are an AI assistant specialized in processing, analyzing, and synthesizing information from multiple sources to answer user questions, generate insights, and prepare detailed reports. Your core capabilities include financial data interpretation, knowledge retrieval, and multi-source analysis, with a focus on supporting operational and strategic decision-making for Golden Section’s portfolio companies.
In performing your duties, you must ensure that all input data undergoes rigorous verification for authenticity, accuracy, and completeness before any analysis is conducted. This includes implementing protocols to validate financial figures, legal terms, and other critical information for alignment with established norms and credible sources. Your process must involve:
  • Cross-referencing information from diverse, credible sources to detect and mitigate false, exaggerated, or incomplete data.
  • Assessing the reliability of each source, prioritizing primary sources where available.
  • Identifying and resolving inconsistencies, ambiguities, or potential misinformation through systematic checks.
  • Ensuring all risk assessments and conclusions are based solely on validated and accurate inputs to maintain the integrity of your outputs.
Data Authenticity and Completeness Verification
  • Scrutinize contextual cues (e.g., “Context: Section: Payback Period:”) to ensure alignment with expected data types and structures.
  • Check for numerical or factual inconsistencies, such as typos (e.g., “11975” instead of “1975”), exaggerations (e.g., “\$22.0M” without supporting context), or missing critical information.
  • Validate that all referenced data points are present, logically consistent, and contextually appropriate.
  • Flag and document any anomalies for further investigation before proceeding with classification or analysis.
Context Interpretation and Parsing Guidelines
  • Carefully interpret and utilize contextual cues, especially in nested or ambiguous contexts (e.g., “Context: Section: Name & Headquarters:”).
  • Accurately parse section headers and contextual clues to prevent misclassification or incomplete analysis in multi-section reports.
  • Anchor analysis to the document’s structure by adhering to hierarchical or sequential organization.
  • Resolve discrepancies in contextual labeling or structure to maintain coherence.
Structured Classification and Risk Prioritization Framework
  1. (1)
    Factual Reporting and Descriptive Analysis: Present verified information such as corporate history, operational metrics, and financial data neutrally, before transitioning to evaluative content.
  2. (2)
    Business Analysis: Evaluate performance, market positioning, and strategic initiatives; assess risks by severity, likelihood, and propose contextualized mitigation.
  3. (3)
    Legal Risk Analysis: Examine compliance, regulatory, and contractual risks; assess impact and propose mitigation actions aligned with legal context.
  4. (4)
    Cross-Domain Analysis: For overlapping elements, classify by primary context and document dual-category cases with rationale.
Validation Mechanisms for Cross-References
  • Distinguish between source types (e.g., governance vs. identity records).
  • For each statement, explicitly identify the source type and ensure contextual alignment.
Empirical Observation. After applying MAPGD, we observe a notable increase in the comprehensiveness of the generated financial analysis reports. Specifically, the average report length increased from 1616 words under the original system prompt to 1965 words with the optimized prompt. This indicates that the optimized prompt encourages the model to produce more detailed and contextually grounded content. While the per-report output length becomes longer, the improved accuracy and completeness reduce the need for repeated generations or manual corrections, thereby lowering the overall token consumption in practical workflows. These findings further support that MAPGD enhances the robustness and effectiveness of system prompts in real-world generation tasks.
Appendix E Agent Specialization for Different Tasks
To adapt MAPGD to different domains, we define specialized agent roles tailored to the unique requirements of each task. Below are the agent configurations used for the classification, mathematical reasoning, and financial analysis tasks in our experiments.
E.1. Agent Roles for Classification Tasks
For general classification tasks, the agents focus on the core components of a good prompt: instructions, examples, format, and style.
instruction_specialist:
Analyzes and improves task instructions, ensuring clarity, completeness, and executability.
example_curator:
Focuses on selecting representative, diverse examples and ensuring consistent formatting.
format_designer:
Designs clear output templates and structured formats for better model understanding.
style_optimizer:
Optimizes language expression for professionalism and task-specific adaptation.
E.2. Agent Roles for Mathematical Reasoning Tasks
For mathematical reasoning, the agent roles are adapted to focus on logical decomposition, calculation, and problem interpretation.
reasoning_specialist:
Specializes in enhancing mathematical reasoning processes, ensuring clear step-by-step logic and proper problem decomposition.
calculation_optimizer:
Focuses on improving calculation accuracy, suggesting better computational methods, and ensuring numerical correctness.
problem_interpreter:
Specializes in interpreting math word problems, extracting relevant information, and identifying mathematical relationships.
solution_formatter:
Optimizes mathematical solution presentation, ensures clear formatting and proper answer notation (e.g., #### [final_answer]).
E.3. Agent Roles for Financial Report Analysis
For the complex task of business and financial analysis, agents are given highly specialized roles focusing on data verification, structured report generation, and professional tone.
instruction_specialist:
Crafts clear, executable instructions for generating truthful, structured reports and rejecting false or exaggerated inputs. Instructions must include steps to verify input data authenticity (e.g., checking financial metrics for realism).
example_curator:
Selects and formats diverse examples that demonstrate truthful report generation and false data rejection, including both positive (realistic) and negative (exaggerated) samples.
format_designer:
Specializes in designing clear, structured output templates for reports (e.g., using <Business Analysis>, <Financial Report> tags) and error messages (e.g., <Error>).
style_optimizer:
Optimizes language style and tone for generating professional reports in business, financial, and legal contexts, while ensuring error messages are direct and unambiguous.
Appendix F Example Prompts Before and After Optimization
This appendix provides the initial prompts and the final optimized prompts generated by MAPGD for each of the benchmark datasets used in our experiments.
F.1. LIAR Dataset
Initial Prompt
# Task
Determine whether the Statement is a lie (Yes) or not (No)
based on the Context and other information.
# Output format
Answer Yes or No as labels
# Prediction
Text: {text}
label:
Optimized Prompt
# Task
Your task is to evaluate the given Statement in the context
of the provided information. If the Statement is found to be
intentionally false or misleading, it should be classified
as a lie. In this case, your response should be ’Yes’.
If the Statement is true or unintentionally false due to a
mistake or lack of knowledge, it should not be classified
as a lie. In this case, your response should be ’No’.
# Output format
Provide your answer as ’Yes’ or ’No’ labels. ’Yes’ indicates
that the statement is a lie, while ’No’ indicates that the
statement is not a lie.
# Prediction
Text: {text}
label:
F.2. Ethos Dataset
Initial Prompt
# Task
Is the following text hate speech?
# Output format
Answer Yes or No as labels
# Prediction
Text: {text}
Label:
Optimized Prompt
# Task
Your task is to analyze the following text and determine if
it contains hate speech. For the purpose of this task, hate
speech is defined as any speech, gesture, conduct, writing,
or display that may incite violence or prejudicial actions
against or by a protected individual or group, or because it
disparages or intimidates a protected individual or group.
Please base your judgment on the explicit content of the text,
rather than personal interpretations or assumptions.
# Output format
Provide your answer in a clear format as either
’Contains Hate Speech’ or ’Does Not Contain Hate Speech’
# Prediction
Text: {text}
Label:
F.3. Jailbreak Dataset
Initial Prompt
# Task
Detect if the message is a jailbreak attack,
i.e. an attempt by a user to break through an
AI system’s protections
# Output format
Answer Yes or No as labels
# Prediction
Text: {text}
Label:
Optimized Prompt
# Task
Classify the message as one of the following labels based on
its potential to be a jailbreak attack, i.e., an attempt to
bypass an AI system’s protections:
- "OvertAttack": The message contains clear, direct attempts
  to break through system safeguards.
- "SuspiciousContext": The message shows indirect or
  contextual cues that may indicate a multi-stage or
  sophisticated attack attempt.
- "Benign": The message poses no apparent risk and does not
  attempt to circumvent protections.
# Output format
Select and output exactly one label from the options above.
# Prediction
Text: {text}
Label:
F.4. GSM8K Dataset
Initial Prompt
# Task
Solve the math word problem step by step.
# Instructions
1. Read the problem carefully
2. Identify what needs to be calculated
3. Show your work step by step
4. Provide the final numerical answer
# Output Format
Show your reasoning process and end with: #### [final_answer]
# Problem
{text}
Optimized Prompt
# Task
Solve the math word problem step by step.
# Instructions
1. Read the problem carefully
2. Identify what needs to be calculated
3. If there are ambiguous situations in the problem, make
   reasonable assumptions and state them clearly
4. Show your work step by step
5. Provide the final numerical answer
# Output Format
Show your reasoning process, including any assumptions made,
and end with: #### [final_answer]
# Problem
{text}
F.5. AQUARAT Dataset
Initial Prompt
# Task
Solve the math word problem and choose the correct answer
from the given options.
# Instructions
1. Read the problem carefully
2. Analyze each option
3. Show your reasoning step by step
4. Select the correct answer (A, B, C, D, or E)
# Output Format
Show your reasoning and end with: Answer: [LETTER]
(For example: "Answer: A" or "Answer: B")
# Problem
{text}
# Options
{options}
Optimized Prompt
# Task
Solve the provided problem by systematically classifying its
type, analyzing all constraints, applying formal mathematical
reasoning, and verifying the solution against all conditions
before selecting the correct answer.
# Instructions
Adhere strictly to the following structured framework:
1. **Problem Classification:**
   * Explicitly identify the core problem domain (e.g.,
     "Combinatorial Selection," "Constraint Satisfaction,"
     "Partnership Profit-Sharing," "Work Rate").
   * Justify the classification with a brief rationale
     based on the problem statement.
2. **Constraint Analysis:**
   * Enumerate all explicit constraints from the problem text.
   * Infer and list any implicit constraints or logical
     dependencies.
   * Categorize each constraint (e.g., Hard/Must-Fulfill,
     Soft/Optimization) and specify their logical
     relationships (e.g., AND, OR).
3. **Mathematical Formalization & Numerical Context
     Resolution:**
   * Identify and resolve any ambiguous numerical references
     (e.g., "monthly" vs. "annual", "together" vs.
     "individual").
   * Infer and incorporate any implicit numerical variables
     (e.g., initial quantities, unstated rates, starting
     points).
   * Normalize all units to a consistent basis for calculation.
   * Define all variables, sets, and parameters using clear
     notation (e.g., C_total, P_A, S_eligible, nCr).
   * Structure the problem using appropriate formalisms:
     timelines, sets, equations, or logical statements.
   * For combinatorial problems, explicitly state
     the formula used.
4. **Solution Execution:**
   * Perform all calculations step-by-step, showing
     substitutions and operations.
   * For combinatorial scenarios, enumerate valid
     combinations or calculate cardinalities.
   * Derive the target value (e.g., profit share, number of
     committees, optimal value).
5. **Verification & Solution Finalization:**
   * Systematically verify that the proposed solution
     satisfies every constraint from Step 2.
   * For combinatorial answers, confirm counts against
     constraints using direct checks or complementary
     counting.
   * Match the final, verified result to the provided
     options to select the correct answer.
# Output Format
Your final output must follow this structure precisely:
**Problem Classification:**
[Your classification and rationale]
**Constraint
F.6. SVAMP Dataset
Initial Prompt
# Task
Solve the math word problem step by step.
# Instructions
1. Read the problem carefully
2. Identify what needs to be calculated
3. Show your work step by step
4. Provide the final numerical answer
# Output Format
Show your reasoning process and end with: #### [final_answer]
# Problem
{text}
Optimized Prompt
F.7. SVAMP Dataset
Optimized Prompt
# Task
Solve the math word problem through structured, hierarchical
reasoning. Your primary goal is to correctly parse and compute
multi-step problems involving distinct object categories and
their quantitative relationships.
# Instructions
1. **Parse and Categorize:**
   Read the problem and question carefully. Identify and list
   all distinct object types or categories (e.g., ’packages
   of gum’, ’boxes of candy’). For each category, extract:
   - The number of units (e.g., 5 packages).
   - The quantity per unit, if specified (e.g., 8 pieces
     per package).
2. **Structure the Solution Hierarchically:**
   - **Step 1: Calculate Intra-Category Totals.**
     For each identified category, calculate its total quantity.
     If a unit rate is given, perform the multiplication
     (e.g., ‘5 packages * 8 pieces/package = 40 pieces of gum‘).
     If no unit rate exists, use the given quantity directly.
   - **Step 2: Perform Inter-Category Operations.**
     Using the totals from Step 1, now perform the final operation
     as required by the question (e.g., sum all category totals
     for a grand total, or find the difference between two
     category totals for a comparison).
3. **Show Your Work:**
   Present your reasoning clearly, reflecting this two-step
   hierarchical process. First, show all calculations within
   categories. Then, show the final combination of these
   category totals.
4. **Final Answer:**
   Box your final numerical result.
# Output Format
Reasoning: [Your step-by-step reasoning here]
#### [final_answer]
# Problem
{text}
# Question
{question}

### v3

@@
 MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization
-Abstract.
-Prompt engineering is essential for fully leveraging large language models (LLMs), yet existing optimization methods often follow a single trajectory, leading to limited adaptability, gradient conflicts, and high computational overhead.
-We propose MAPGD (Multi-Agent Prompt Gradient Descent), a novel framework that reconceptualizes prompt optimization as a collaborative multi-agent process.
-MAPGD incorporates specialized agents focusing on orthogonal refinement dimensions—such as instruction clarity, example selection, format structuring, and stylistic adaptation—and coordinates their contributions through semantic gradient embedding, conflict detection, and fusion.
-To further enhance robustness and stability, MAPGD proposes two new mechanisms: Hypersphere-Constrained Gradient Clustering (HCGC), which enforces angular margin constraints to ensure intra-cluster compactness and inter-cluster separation, and Channel-Adaptive Agent Weighting (CAAW), which dynamically reweights agent contributions based on validation performance.
-Experiments on classification and reasoning benchmarks demonstrate that MAPGD consistently outperforms single-agent and random baselines in both accuracy and efficiency.
-Ablation studies validate the effectiveness of gradient fusion, agent specialization, and conflict resolution.
-Together, these contributions establish MAPGD as a unified, gradient-inspired, and interpretable framework for robust prompt optimization with theoretical convergence guarantees.
-Key words and phrases:
-Prompt Optimization, Multi-Agent Systems, Large Language Models, Gradient Descent
-{CCSXML}
+Abstract
+Prompt engineering is crucial for fully leveraging large language models, yet existing prompt optimization methods often rely on a single refinement trajectory, leading to instability, conflicting update signals, and inefficient use of query budgets.
+We propose Multi-Agent Prompt Gradient Descent (MAPGD), a gradient-inspired framework that coordinates multiple complementary prompt editing signals for robust and interpretable optimization.
+MAPGD decomposes prompt refinement into orthogonal dimensions (e.g., instruction clarity, example selection, format, and style) and aggregates textual pseudo-gradients via semantic embedding, conflict-aware clustering, and adaptive fusion.
+To improve robustness, we introduce HCGC, which enforces angular separation between conflicting directions, and CAAW, which dynamically calibrates editing contributions based on validation feedback.
+Experiments on diverse classification and reasoning benchmarks show that MAPGD achieves consistent gains in accuracy over strong baselines, while ablation results indicate that coordinated gradient fusion and adaptive weighting are critical to stable optimization.
+Together, these findings suggest that MAPGD offers a robust and interpretable framework for automatic prompt optimization.
+MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization
 […]
-1. Introduction
-Figure 1. Overview of the Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization.
-[…]
-Overview of the Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization.
-Large Language Models (LLMs), trained on vast webscale corpora, have recently demonstrated remarkable generalization capabilities across diverse natural language processing (NLP) tasks Achiam et al.
+1 Introduction
+Large Language Models (LLMs), trained on vast web-scale corpora, have demonstrated remarkable generalization capabilities across a wide range of natural language processing (NLP) tasks, including classification, reasoning, and generation Achiam et al.
 (2023); Bubeck et al.
 (2023).
-A central factor influencing their performance is the design of prompts, which serve as the primary interface for guiding model behavior.
-However, prompt construction remains largely a manual, trial-and-error process that demands substantial human labor Jiang et al.
+A critical factor underlying this performance is the design of prompts, which serve as the primary interface for steering model behavior.
+Despite their importance, prompts are still predominantly crafted through manual trial-and-error, requiring substantial human effort Jiang et al.
 (2022) and domain expertise Reynolds and McDonell (2021); Zamfirescu-Pereira et al.
 (2023).
-This inefficiency highlights the urgent need for automatic or semi-automatic strategies to systematically generate or refine prompts.
-Such approaches can reduce manual burden, improve task performance, and provide interpretable optimization pathways.
-Recent research has explored this problem through two primary directions.
-One line of work introduces auxiliary models or differentiable prompt embeddings to approximate gradient-based optimization Qin and Eisner (2021); Deng et al.
+This reliance on manual design limits scalability and motivates the development of automatic or semi-automatic prompt optimization methods that are both effective and interpretable.
+Existing approaches to prompt optimization largely fall into two categories.
+The first line of work approximates gradient-based optimization through auxiliary models or differentiable prompt representations Qin and Eisner (2021); Deng et al.
 (2022).
-Yet, these methods often assume access to the internal states or parameters of LLMs Shin et al.
+While theoretically appealing, these methods often assume access to internal model states or parameters Shin et al.
 (2020); Lester et al.
-(2021), which is impractical in real-world API-only scenarios.
-Another line of work employs reinforcement learning or feedback-driven editing to discretely manipulate prompts Zhang et al.
+(2021), making them impractical in real-world, API-only settings.
+The second line of work formulates prompt optimization as a discrete search or reinforcement learning problem, using feedback-driven edits or Monte Carlo exploration Zhang et al.
 (2022); Zhou et al.
 (2022).
-While effective in some cases, these algorithms may produce incoherent prompts, incur high computational costs, or rely on unguided Monte Carlo search over the semantic space.
-Collectively, these limitations highlight the need for more robust, interpretable, and scalable frameworks for prompt optimization.
-Addressing this gap, Pryzant et al.
+Although effective in certain cases, these methods can suffer from high computational cost, unstable optimization trajectories, and limited interpretability.
+Collectively, these limitations highlight the need for more robust, scalable, and transparent prompt optimization frameworks.
+To bridge this gap, Pryzant et al.
 Pryzant et al.
-(2023) proposed ProTeGi, a nonparametric approach that mirrors gradient descent in textual space.
-ProTeGi generates natural language “gradients” from model errors and iteratively edits prompts in semantically opposite directions.
-To balance exploration and exploitation, it further integrates beam search and bandit-based candidate selection, achieving significant gains over earlier baselines.
-Nevertheless, ProTeGi operates under a single-agent paradigm, which restricts the diversity of optimization signals and leaves unresolved conflicts among competing refinement directions.
-To overcome these limitations, we propose MAPGD (Multi-Agent Prompt Gradient Descent), a novel framework that reconceptualizes prompt optimization as a collaborative multi-agent process(Figure 1).
-MAPGD introduces a set of specialized agents, each dedicated to distinct refinement dimensions—such as instruction clarity, example selection, format structuring, and stylistic adaptation.
-To reconcile heterogeneous improvement signals, MAPGD employs a semantic gradient coordinator that embeds textual gradients into a shared vector space, enabling conflict detection, clustering, and fusion.
-In addition, we propose two novel mechanisms to further enhance robustness and stability: Hypersphere-Constrained Gradient Clustering (HCGC), which introduces angular margin constraints on the unit hypersphere to ensure compact intra-cluster coherence and inter-cluster separation, and Channel-Adaptive Agent Weighting (CAAW), which dynamically reweights agent contributions according to validation performance, amplifying effective optimization channels while suppressing noisy ones.
-Together, these innovations endow MAPGD with improved robustness, stability, and interpretability, enabling it to consistently outperform single-agent baselines such as ProTeGi across diverse optimization scenarios.
-Beyond empirical improvements, MAPGD also provides a theoretical foundation for convergence.
-Under mild stochastic approximation assumptions, we prove that the multi-agent gradient coordination and fusion mechanisms preserve the convergence properties of classical gradient descent, yielding almost sure convergence to a local optimum at a sublinear rate of $\mathcal{O}(1/\sqrt{T})$, where $T$ denotes the number of iterations.
-This result establishes that, despite operating in a discrete, textual prompt space and coordinating multiple heterogeneous agents, MAPGD maintains stable and predictable optimization dynamics, bridging practical prompt refinement with rigorous theoretical guarantees.
-Figure 2.
-Workflow of MAPGD.
-Starting from an initial prompt $p$, specialized agents propose diverse pseudo-gradients $\{g,g^{\prime},g^{\prime\prime},g^{\prime\prime\prime},\ldots\}$ (top).
-These gradients are semantically embedded and projected onto the unit hypersphere, then clustered with angular-margin constraints to form coherent groups (middle).
-Clustered directions are fused via LLM-driven synthesis into composite gradients $G$, which generate refined candidate prompts $\{P_{1},P_{2},\ldots,P_{n}\}$ (bottom).
-The best candidates are iteratively selected and fed back to continue the prompt optimization loop.
-[…]
-Workflow of MAPGD.
-Starting from an initial prompt $p$, specialized agents propose diverse pseudo-gradients $\{g,g^{\prime},g^{\prime\prime},g^{\prime\prime\prime}$…\} (top).
-These gradients are semantically embedded and projected onto the unit hypersphere, then clustered with angular-margin constraints to form coherent groups (middle).
-Clustered directions are fused via LLM-driven synthesis into composite gradients $G$, which generate refined candidate prompts $\{P_{1},P_{2},\ldots,P_{n}\}$ (bottom).
-The best candidates are iteratively selected and fed back to continue the prompt optimization loop.
-2. Related Work
+(2023) proposed ProTeGi, a non-parametric framework that reinterprets prompt optimization as gradient descent in natural language space.
+ProTeGi generates textual pseudo-gradients from model errors and iteratively refines prompts using beam search and bandit-based selection, achieving strong empirical performance.
+However, ProTeGi relies on a single-agent optimization framework that aggregates heterogeneous refinement signals into a unified trajectory.
+Such entanglement may give rise to implicit objective conflicts, thereby impairing optimization stability and convergence quality.
+A natural extension is to decompose prompt optimization into multiple specialized perspectives.
+Multi-agent formulations enable parallel exploration of heterogeneous refinement dimensions, increasing signal diversity and reducing premature convergence.
+However, introducing multiple agents also makes semantic conflicts inevitable: different agents may propose mutually incompatible updates, such as simultaneously expanding illustrative examples while enforcing concise instructions.
+Without explicit coordination mechanisms, such conflicts can destabilize optimization and undermine the benefits of multi-agent collaboration.
+In this work, we propose MAPGD, a unified framework that addresses these challenges by integrating multi-agent specialization with principled coordination mechanisms (Figure 1).
+As illustrated in Figure 1(a), MAPGD decomposes prompt optimization into multiple orthogonal refinement dimensions and assigns each to a specialized agent.
+Each agent independently analyzes model errors and produces textual pseudo-gradients that capture targeted improvement directions.
+However, pseudo-gradients produced by different agents may encode semantically conflicting refinement intents.
+To explicitly disentangle such conflicts, MAPGD proposes Hypersphere-Constrained Gradient Clustering (HCGC), which embeds agent-generated gradients into a shared semantic space and separates incompatible optimization directions via angularly constrained clustering.
+This design promotes semantic consistency within clusters while preventing destructive interference across conflicting refinement signals.
+While HCGC addresses conflicts at the directional level, agent contributions within the same semantic cluster may still vary in reliability.
+MAPGD therefore incorporates CAAW, which dynamically reweights agent channels based on validation performance.
+By emphasizing consistently effective signals and attenuating noisy or unstable ones, CAAW enables robust aggregation of non-conflicting gradients, leading to stable and effective prompt refinement.
+In summary, our contributions are three-fold.
+(1) We propose MAPGD, a multi-agent pseudo-gradient framework for stable and interpretable prompt optimization.
+(2) We introduce HCGC and CAAW to disentangle conflicting refinement signals and adaptively aggregate reliable agent contributions.
+(3) We establish theoretical convergence guarantees, showing that MAPGD preserves classical gradient descent behavior with almost sure convergence to a local optimum at rate $\mathcal{O}(1/\sqrt{T})$.
+Figure 1: Overview of MAPGD.
+MAPGD performs prompt optimization through collaborative multi-agent pseudo-gradients.
+Agent-generated gradients are embedded and clustered via HCGC to resolve semantic conflicts, and adaptively weighted by CAAW according to validation performance.
+The resulting fused gradients iteratively refine the prompt through candidate generation and selection.
+2 Related Work
 Prompt Learning and Optimization.
-Early efforts relied on manual engineering, which lacks scalability.
-Automated methods include reinforcement learning Deng et al.
+Early prompt engineering relied on manual design, which is labor-intensive and difficult to scale.
+Automated approaches have since been explored, including reinforcement learning Deng et al.
 (2022), evolutionary search Fernando et al.
@@
 (2025).
-Continuous prompt tuning (e.g., prefix tuning Li and Liang (2021), soft prompts Lester et al.
-(2021)) optimizes embeddings but sacrifices interpretability.
-MAPGD instead pursues interpretable natural language optimization with structured feedback.
+Continuous prompt tuning methods, such as prefix tuning Li and Liang (2021) and soft prompts Lester et al.
+(2021), optimize prompts in embedding space but often sacrifice interpretability and require access to model internals.
+MAPGD operates in natural language space, preserving interpretability while enabling feedback-driven optimization.
+Gradient-Inspired Prompt Optimization.
+Recent work has framed prompt optimization as a gradient-like process in discrete language space.
+ProTeGi Pryzant et al.
+(2023) generates natural language pseudo-gradients from model errors and iteratively refines prompts via beam search and bandit-based selection.
+However, as a single-agent method, ProTeGi entangles heterogeneous refinement objectives and lacks explicit mechanisms for resolving conflicting optimization signals, which can lead to unstable or suboptimal convergence.
+This limitation motivates approaches that can disentangle and coordinate diverse refinement directions.
 Multi-Agent Collaboration.
-Multi-agent systems are well-studied in RL, distributed AI, and game theory.
+Multi-agent formulations enable parallel exploration of diverse perspectives and have been widely studied in reinforcement learning and distributed AI.
 In NLP, multi-agent debate Liang et al.
 (2023); Du et al.
-(2023), cooperative reasoning, and collaborative generation Hong et al.
-(2024) show that role specialization outperforms monolithic optimization.
-MAPGD applies this principle by assigning agents to distinct prompt dimensions and coordinating them via gradient fusion.
-Gradient-Inspired Optimization.
-ProTeGi Pryzant et al.
-(2023) and similar approaches approximate prompt gradients through self-feedback.
-While effective, single-agent descent struggles with limited signal diversity and inherent conflict resolution challenges Xiang et al.
+(2023) and collaborative generation Hong et al.
+(2024) demonstrate the benefits of role specialization.
+However, increased diversity makes semantic conflicts inevitable, as different agents may propose incompatible updates.
+MAPGD adopts a multi-agent design with specialized roles, while explicitly addressing the resulting coordination and conflict resolution challenges.
+Geometric and Adaptive Coordination.
+Resolving semantic conflicts requires structured inductive biases beyond heuristic fusion.
+In representation learning, contrastive objectives and angular margin constraints enforce intra-cluster compactness and inter-cluster separation Wang and Isola (2020); Liu et al.
+(2017); Wang et al.
+(2018); Deng et al.
 (2019).
-A monolithic agent can generate self-contradictory pseudo-gradients, as different refinement dimensions are not explicitly disentangled.
-For instance, to correct a misclassification, a single agent might propose a gradient to add more illustrative examples, while an implicit counter-gradient aims to make the instructions more concise.
-This creates a direct conflict in prompt length and structure, potentially leading to suboptimal or oscillating optimization trajectories.
-MAPGD extends this paradigm by embedding gradients semantically, explicitly detecting such conflicts, and fusing them coherently through LLM-driven synthesis.
-Geometric and Adaptive Mechanisms.
-Contrastive learning and angular margin constraints are effective in representation learning.
-Wang and Isola formalize how contrastive losses enforce alignment and uniformity, providing a theoretical lens Wang and Isola (2020).
-This principle of hyperspherical separation to enhance discriminability spans domains.
-In face recognition, margin-based softmax losses such as SphereFace Liu et al.
-(2017), CosFace Wang et al.
-(2018), and ArcFace Deng et al.
-(2019) enforce angular margins between identities.
 Chen et al.
 Chen et al.
-(2021) introduce a distance polarization regularizer for large-margin contrastive learning, pushing inter-cluster samples apart while keeping intra-cluster representations tight.
-Inspired by these advances, our Hypersphere-Constrained Gradient Clustering (HCGC) enforces intra-cluster compactness and inter-cluster separation for semantic gradients, complemented by Channel-Adaptive Agent Weighting (CAAW) adjusting agent contributions.
-Together, these modules enhance multi-agent gradient descent with geometry-aware clustering and adaptive coordination.
-In summary, MAPGD unifies prompt learning, multi-agent collaboration, gradient descent, and geometry-inspired optimization into a scalable and interpretable framework.
-3. Methodology
-3.1. Framework Overview
-MAPGD formulates prompt optimization as a hybrid discrete–continuous gradient descent process in the natural language space.
-Unlike continuous embedding–based methods such as prefix or soft prompt tuning Li and Liang (2021); Lester et al.
-(2021), MAPGD explicitly manipulates interpretable textual prompts while leveraging gradient-inspired feedback for iterative refinement (Figure 2).
-At each iteration, specialized agents analyze misclassified or suboptimal examples and produce pseudo-gradients in natural language.
-These gradients are semantically embedded, clustered, and adaptively fused to mitigate conflicts and ensure coherent optimization.
-The resulting fused gradients are applied to generate successor prompts, which are then evaluated and filtered under limited budgets.
-Iterative repetition yields a dynamically evolving prompt trajectory that balances exploration of diverse directions with exploitation of promising refinements.
-Formally, the optimization objective is defined as:
+(2021) further enhance large-margin contrastive learning via distance polarization regularization.
+Inspired by these advances, MAPGD introduces HCGC for geometry-aware gradient coordination, complemented by CAAW to adaptively calibrate agent contributions.
+Concurrent Work.
+Concurrent studies such as MAPRO Zhang et al.
+(2025) and MA-SAPO Seo et al.
+(2025) also explore multi-agent prompt optimization under different assumptions, including explicit interaction structures or external memory components.
+These settings are not directly aligned with our budgeted, task-level optimization protocol.
+A detailed comparison is provided in Appendix E.
+In summary, MAPGD unifies prompt optimization, gradient-inspired refinement, multi-agent collaboration, and geometry-aware coordination into an interpretable framework that explicitly addresses semantic conflicts.
+3 Methodology
+3.1 Framework Overview
+MAPGD formulates prompt optimization as a hybrid discrete–continuous gradient descent process operating directly in natural language space.
+Unlike continuous embedding-based approaches such as prefix tuning or soft prompt optimization Li and Liang (2021); Lester et al.
+(2021), MAPGD preserves interpretability by explicitly manipulating textual prompts while still leveraging gradient-inspired feedback signals for iterative refinement.
+As illustrated in Figure 1, MAPGD frames prompt optimization as a collaborative multi-agent process, in which specialized agents focus on orthogonal refinement dimensions and produce textual pseudo-gradients that guide prompt updates.
+Building on this formulation, MAPGD instantiates the optimization procedure through an iterative coordination loop.
+At each iteration, specialized agents independently examine the current prompt and available task feedback, and generate natural language pseudo-gradients that reflect distinct refinement perspectives.
+These gradients are embedded into a shared semantic space, where potential conflicts are explicitly identified and mitigated via geometry-aware clustering and adaptive fusion.
+The resulting fused gradients guide the generation of successor prompts, which are subsequently evaluated and filtered under a budgeted selection mechanism.
+Repeating this process yields an evolving optimization trajectory that systematically balances exploration across heterogeneous refinement directions with exploitation of effective updates.
+Formally, the prompt optimization objective is defined as
 $F(p)=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\ell(M(x;p),y)\right],$ (1)
-where $\ell(\cdot,\cdot)$ denotes a task-specific loss function. The goal is:
+where $\ell(\cdot,\cdot)$ denotes a task-specific loss function and $M(x;p)$ represents the model output conditioned on prompt $p$. The optimization goal is
 $p^{*}=\arg\min_{p}F(p).$ (2)
-Unlike stochastic gradient descent (SGD), where gradients are continuous, MAPGD constructs textual pseudo-gradients:
+Since gradients in this discrete textual space are not directly accessible, MAPGD constructs natural language pseudo-gradients,
 $\nabla F(p^{(t)})\approx g^{(t)},$ (3)
-which act as semantic analogues of numerical gradients.
-3.2. Specialized Prompt Agents
-Each agent specializes in a distinct dimension of prompt optimization, analogous to orthogonal gradient directions.
-For example: $A_{1}$: instruction clarity ($g_{1}$), $A_{2}$: example selection ($g_{2}$), $A_{3}$: format specification ($g_{3}$), $A_{4}$: stylistic refinement ($g_{4}$), and optionally a generic agent for broad improvements.
-At iteration $t$, the gradient set is:
+which serve as semantic analogues of numerical gradients in classical optimization.
+3.2 Specialized Prompt Agents
+MAPGD employs multiple specialized agents, each responsible for a distinct and approximately orthogonal dimension of prompt refinement.
+Typical agent roles include instruction clarity, example selection, output formatting, and stylistic refinement, optionally complemented by a generic agent for broad exploratory updates.
+At iteration $t$, the resulting pseudo-gradient set is
 $G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}.$
-This decomposition supports parallel exploration of diverse refinements, alleviating local minima issues inherent in single-agent approaches.
-3.3. Hypersphere-Constrained Gradient Clustering
-Inspired by hyperspherical contrastive learning Wang et al.
+By decomposing prompt optimization into specialized perspectives, MAPGD enables parallel exploration of heterogeneous refinement directions, alleviating the limited signal diversity and premature convergence commonly observed in single-agent optimization.
+3.3 Hypersphere-Constrained Gradient Clustering
+Inspired by hyperspherical metric learning Wang et al.
 (2018); Liu et al.
-(2017), our HCGC mechanism is designed to enhance the semantic consistency of multi-agent gradients by enforcing intra-cluster compactness and inter-cluster separation in a normalized hypersphere space.
-This approach addresses the challenge of conflicting pseudo-gradients generated by heterogeneous agents, ensuring that semantically coherent optimization directions are emphasized while contradictory ones are disentangled.
-The overall process and the geometric intuition of our method are illustrated in Figure 3.
-Gradient Embedding on Hypersphere.
-Given a set of agent-generated gradients $G^{(t)}=\{g_{1}^{(t)},g_{2}^{(t)},\dots,g_{K}^{(t)}\}$, each gradient is encoded into a d-dimensional vector $v_{k}^{(t)}=\phi(g_{k}^{(t)})$, where $\phi$ denotes a pre-trained language encoder.
-For this purpose, we utilize a Sentence Transformer model, such as all-MiniLM-L6-v2, which provides compact and semantically informative embeddings.
-To eliminate scale variance, we normalize each embedding onto the unit hypersphere:
+(2017), HCGC resolves semantic conflicts among heterogeneous pseudo-gradients by enforcing geometry-aware separation on a normalized hypersphere.
+Instead of assuming all agent proposals are mutually compatible, HCGC explicitly detects semantic disagreement and clusters gradients into coherent refinement directions before fusion.
+Gradient Embedding and Conflict Signal.
+Given agent-generated pseudo-gradients $G^{(t)}$, each $g_{k}^{(t)}$ is encoded into a $d$-dimensional semantic vector $v_{k}^{(t)}=\phi(g_{k}^{(t)})$ and normalized onto the unit hypersphere:
 $\hat{v}_{k}^{(t)}=\frac{v_{k}^{(t)}}{\|v_{k}^{(t)}\|},\quad\hat{v}_{k}^{(t)}\in\mathbb{S}^{d-1}.$ (4)
-The similarity between two gradients is then computed as the cosine similarity.
-Gradients are considered to be in conflict if their angular distance exceeds a predefined threshold, $\theta_{conflict}$:
+Semantic similarity is measured by cosine similarity,
 $\mathrm{sim}(\hat{v}_{i},\hat{v}_{j})=\hat{v}_{i}^{\top}\hat{v}_{j}=\cos(\Delta(\hat{v}_{i},\hat{v}_{j})),$ (5)
 where $\Delta(\cdot,\cdot)$ denotes the angular distance on the hypersphere.
-Figure 4 illustrates a concrete example in the sentiment classification task, where different agents generate multiple pseudo-gradients.
-While some gradients form semantically coherent clusters (intra-cluster compactness), others introduce conflicts that must be separated through angular margin constraints.
+To make “conflict” explicit and measurable, we define two gradients as being in semantic conflict if their similarity falls below a threshold:
+$\mathrm{conflict}(\hat{v}_{i},\hat{v}_{j})\;\Leftrightarrow\;\hat{v}_{i}^{\top}\hat{v}_{j}<\theta_{\text{conflict}}.$ (6)
+Intuitively, low cosine similarity indicates substantially different (and potentially incompatible) refinement intents, motivating explicit separation in clustering.
+Clustering and Adaptive Cluster Cardinality.
+We perform cosine K-means on the hypersphere to group semantically coherent directions. Rather than fixing the number of clusters, we adopt a bounded adaptive strategy:
+$K^{(t)}=\min\big(|G^{(t)}|,\;K_{\max}\big),$ (7)
+where $|G^{(t)}|$ is the number of gradients at iteration $t$ and $K_{\max}$ is an upper bound.
+This prevents over-fragmentation when gradients are few, while allowing multiple distinct directions when agent proposals are diverse.
 Angular Margin Constraint.
-To mitigate conflicts and ensure clear separation between semantic clusters, HCGC enforces an angular margin constraint on the unit hypersphere.
-Let $\hat{v}_{k}$ denote a normalized gradient vector, $u_{i}$ the centroid of its assigned cluster $i$, and $u_{j}$ the centroid of any other cluster $j$.
-We define the corresponding angles as
-$\alpha=\Delta(\hat{v}_{k},u_{i}),\quad\beta=\Delta(\hat{v}_{k},u_{j}).$
-A standard clustering assignment only requires $\alpha<\beta$. However, to enforce stronger separation, we introduce a margin scaling factor $n\geq 1$:
-$n\cdot\alpha<\beta.$
-Since $\cos(\theta)$ is monotonically decreasing for $\theta\in[0,\pi]$, this condition is equivalently expressed in terms of cosine similarity:
-$\cos(n\cdot\alpha)>\cos(\beta),\quad n\geq 1.$ (6)
-This constraint ensures that $\hat{v}_{k}$ is not only closer to its own centroid $u_{i}$, but also significantly more aligned with it than with any other centroid $u_{j}$.
-Geometrically, as illustrated in Figure 3 (c), this confines $\hat{v}_{k}$ to a scaled angular neighborhood around $u_{i}$, effectively creating a “cone of acceptance” narrower than the standard Voronoi region on the hypersphere.
-By enforcing this margin, HCGC simultaneously enhances inter-cluster separation and intra-cluster compactness, yielding more stable and reliable gradient fusion.
-During clustering, any gradient $\hat{v}_{k}$ violating this condition with respect to some $u_{j}$ is considered for reassignment to a cluster where the margin holds; otherwise, it remains in its original cluster.
-Figure 3. Details of Hypersphere-Constrained Gradient Clustering
+Initial clustering may still yield weakly separated clusters when conflicting gradients coexist.
+HCGC therefore refines assignments with an angular margin constraint.
+Let $u_{i}$ be the centroid of $\hat{v}_{k}$’s assigned cluster and $u_{j}$ any other centroid.
+Define $\alpha=\Delta(\hat{v}_{k},u_{i})$ and $\beta=\Delta(\hat{v}_{k},u_{j})$.
+While standard assignment requires $\alpha<\beta$, HCGC strengthens it by a margin scale $n\geq 1$:
+$n\cdot\alpha<\beta,$ (8)
+equivalently,
+$\cos(n\cdot\alpha)>\cos(\beta).$ (9)
+Gradients that violate the margin are reassigned to clusters where the constraint holds, yielding more compact and better-separated semantic groups.
+Fusion of Clustered Gradients.
+Finally, gradients within each cluster are fused into a single representative refinement direction via LLM-guided synthesis.
+To avoid redundant updates, we apply a diversity filter to discard fused gradients overly similar to those already selected, controlled by $\theta_{\text{diversity}}$.
+Figure 2: Effect of hypersphere-constrained angular margin in HCGC.
+(a) Clustering without angular margin constraints may entangle semantically conflicting pseudo-gradients, leading to ambiguous cluster boundaries.
+(b) Enforcing an angular margin on the unit hypersphere improves intra-cluster compactness and inter-cluster separation, effectively isolating conflicting refinement directions.
+3.4 Channel-Adaptive Agent Weighting
+Although HCGC resolves structural conflicts between semantic gradient directions, gradients within the same cluster may still vary in reliability due to heterogeneous agent expertise and historical effectiveness.
+To account for this variability, we introduce CAAW, a reliability-aware coordination mechanism that adaptively calibrates agent contributions during gradient fusion.
+CAAW is inspired by channel-wise recalibration mechanisms in neural representations Hu et al.
+(2018), but operates at the level of agent-generated textual gradients.
+Adaptive Weight Estimation.
+Let $s_{k}^{(t)}$ denote the validation performance gain attributed to gradient $g_{k}^{(t)}$ at iteration $t$, estimated from historical agent performance statistics.
+CAAW converts these scalar signals into normalized reliability weights via a softmax transformation:
+$w_{k}^{(t)}=\frac{\exp(\lambda s_{k}^{(t)})}{\sum_{j\in C}\exp(\lambda s_{j}^{(t)})},$ (10)
+where $C$ denotes the set of gradients within the same semantic cluster and $\lambda$ controls the sharpness of the weighting distribution.
+This formulation assigns higher weights to agents that have consistently produced effective refinements, while down-weighting less reliable channels.
+Weight-Conditioned LLM Fusion.
+Rather than explicitly performing an arithmetic weighted sum over textual gradients, CAAW injects the estimated weights into the LLM-driven fusion operator as conditioning signals.
+Concretely, the fusion operator
+$g_{\text{fused}}^{(t)}=\Psi\big(\{g_{k}^{(t)}\}_{k\in C}\,\big|\,\{w_{k}^{(t)}\}_{k\in C}\big)$ (11)
+receives both the set of clustered gradients and their corresponding reliability weights, which are encoded in the fusion prompt to bias the synthesis process toward more trustworthy agent suggestions.
+In this way, agent weights influence the fused gradient implicitly through prompt-level emphasis, rather than through explicit numerical combination.
+Stability Implications.
+By conditioning gradient fusion on adaptive reliability weights, CAAW suppresses noisy or redundant updates while amplifying consistently effective refinement channels.
+This weight-aware fusion mechanism reduces variance in the optimization trajectory and complements HCGC by addressing intra-cluster reliability differences, thereby promoting stable convergence toward semantically robust prompt candidates.
+3.5 Candidate Generation and Selection
+Fused gradients generate successor prompts $\{p^{\prime}_{1},\dots,p^{\prime}_{n}\}$.
+MAPGD applies lightweight filtering and optional bandit-based evaluation Audibert and Bubeck (2010) to keep the computation budgeted.
+Unlike prior work that depends heavily on bandit selection, MAPGD attains robustness and interpretability mainly via conflict-aware clustering and adaptive weighting across agents.
+Overall algorithm can be found in Algorithm 1.
+Algorithm 1 MAPGD with HCGC and CAAW
+1:Initial prompt $p_{0}$, train data $D_{\text{train}}$, dev data $D_{\text{dev}}$, agents $\{A_{i}\}_{i=1}^{N}$, iterations $R$, beam width $k$
+2:Initialize best prompt $p_{\star}^{(0)}\leftarrow p_{0}$, beam $B_{0}\leftarrow\{p_{0}\}$; initialize each agent prompt $A_{i}.p\leftarrow p_{0}$
+3:for $t=1$ to $R$ do
+4:  $M_{t}\leftarrow\textsc{SampleMiniBatch}(D_{\text{train}},b)$
+5:  $\mathcal{G}_{t}\leftarrow\textsc{GeneAgentGrads}(\{A_{i}\},M_{t})$ $\triangleright$ Alg. 2
+6:  $\tilde{\mathcal{G}}_{t}\leftarrow\textsc{HCGC}(\mathcal{G}_{t})$ $\triangleright$ Alg. 3
+7:  $F_{t}\leftarrow\textsc{CAAW}(\tilde{\mathcal{G}}_{t},D_{\text{dev}})$ $\triangleright$ Alg. 4
+8:  $\mathcal{C}_{t}\leftarrow\textsc{ExpandPrompts}(p_{\star}^{(t-1)},F_{t})$
+9:  $B_{t},p_{\star}^{(t)}\leftarrow\textsc{BanditSelect}(\mathcal{C}_{t},D_{\text{dev}},k)$
 […]
-Details of Hypersphere-Constrained Gradient Clustering: Intra-cluster Compactness and Inter-cluster Repulsion, Original Boundary, Angular Margin
-Optimization Objective.
-Analogous to hyperspherical contrastive learning, the clustering objective combines intra-cluster compactness and inter-cluster repulsion:
-$\mathcal{L}_{\text{HCGC}}=-\log\frac{\exp(\cos(\Delta(\hat{v}_{i},\hat{v}_{j}))/\tau)}{\sum_{k\in\mathcal{N}(i)}\exp(\cos(\Delta(\hat{v}_{i},\hat{v}_{k}))/\tau)},$ (7)
-where $(\hat{v}_{i},\hat{v}_{j})$ is a positive pair from the same cluster, $\mathcal{N}(i)$ denotes negatives from other clusters, and $\tau$ is a temperature factor.
-By optimizing $\mathcal{L}_{\text{HCGC}}$, we obtain semantically well-aligned clusters of gradients, ensuring coherent downstream fusion.
-After clustering, gradients within each semantically coherent group are fused into a single representative gradient.
-This fusion is typically performed by an LLM, guided by the original suggestions.
-To maintain diversity and avoid redundancy among the fused gradients, a semantic diversity filter is applied, discarding suggestions that are too similar to already selected ones, controlled by a diversity threshold $\theta_{diversity}$.
-Figure 4.
-Illustration of agent-generated pseudo-gradients in a sentiment classification task.
-Blue gradients form a coherent cluster under the Instruction Specialist, while the red gradient represents a conflicting update.
-Green gradients from the Example Curator constitute another semantically compact cluster.
-This motivates the need for hypersphere-constrained clustering to enforce intra-cluster compactness and inter-cluster separation.
 […]
-Illustration of agent-generated pseudo-gradients in a sentiment classification task.
-Blue gradients form a coherent cluster under the Instruction Specialist, while the red gradient represents a conflicting update.
-Green gradients from the Example Curator constitute another semantically compact cluster.
-This motivates the need for hypersphere-constrained clustering to enforce intra-cluster compactness and inter-cluster separation.
-3.4. Channel-Adaptive Agent Weighting
-While HCGC ensures structural consistency across gradients, gradients within the same cluster may still differ in reliability due to heterogeneous agent expertise.
-To address this, we propose Channel-Adaptive Agent Weighting (CAAW), which dynamically calibrates agent contributions according to their historical effectiveness, akin to channel-wise recalibration in neural representations Hu et al.
-(2018).
-Adaptive Weight Assignment.
-Let $s_{k}^{(t)}$ denote the validation performance gain associated with gradient $g_{k}^{(t)}$. The adaptive weight of each gradient is computed via a softmax normalization:
-$w_{k}^{(t)}=\frac{\exp(\lambda s_{k}^{(t)})}{\sum_{j\in C}\exp(\lambda s_{j}^{(t)})},$ (8)
-where $C$ is the current cluster, and $\lambda$ controls the sharpness of weighting. This formulation emphasizes historically effective agents while down-weighting less reliable ones.
-Gradient Fusion.
-The final fused gradient is expressed as:
-$g_{\text{fused}}^{(t)}=\Psi\left(\sum_{k\in C}w_{k}^{(t)}g_{k}^{(t)}\right),$ (9)
-where $\Psi(\cdot)$ denotes an LLM-driven synthesis function that integrates multiple textual directions into a coherent refinement.
-Robustness and Stability.
-By recalibrating agent contributions at the channel level, CAAW suppresses redundant or conflicting updates while amplifying stable and discriminative directions.
-Theoretically, this reduces variance in the optimization trajectory and prevents over-amplification of noisy gradients, thereby ensuring convergence to semantically robust prompt candidates.
-3.5. Candidate Generation and Selection
-Fused gradients generate successor prompts $\{p^{\prime}_{1},\dots,p^{\prime}_{n}\}$.
-MAPGD applies lightweight filtering and optional bandit-based evaluation Audibert and Bubeck (2010) to control computational overhead.
-However, unlike prior work that heavily relies on bandit selection, our framework emphasizes geometry-aware clustering (HCGC) and adaptive agent weighting (CAAW) as the primary mechanisms ensuring robustness and interpretability.
-3.6. Theoretical Convergence (Informal)
-We provide a convergence analysis of MAPGD in Appendix C.
-Here we summarize the key assumptions: (i) bounded variance of agent pseudo-gradients, (ii) Lipschitz continuity of the loss function, and (iii) unbiased sampling from bandit-based candidate selection.
-Under these conditions, MAPGD inherits the convergence properties of stochastic gradient descent and achieves almost sure convergence to a local optimum at a sublinear rate of $\mathcal{O}(1/\sqrt{T})$.
-Intuitively, HCGC reduces variance by enforcing geometric separation, CAAW emphasizes historically reliable agents, and bandit selection maintains unbiased exploration, together ensuring stable optimization dynamics.
-4. Experiments
-We design comprehensive experiments to evaluate the effectiveness, efficiency, and robustness of the proposed MAPGD framework.
-To ensure comparability, we reproduce the experimental setup of ProTeGi under identical conditions and extend the evaluation with ablation studies and token consumption analysis.
-In addition, we include a real-world case study on prompt optimization for text generation (see Appendix D), which highlights the applicability of MAPGD beyond controlled benchmarks.
-4.1. Datasets
+12:  end if
+13:end for
+14:return Best prompt over $\{p_{\star}^{(1)},\dots,p_{\star}^{(t)}\}$
+3.6 Theoretical Convergence
+We provide a theoretical analysis of MAPGD in Appendix G.
+Here, we summarize the key assumptions and main implications to clarify how the proposed framework relates to classical stochastic optimization.
+Specifically, our analysis relies on three standard conditions: (i) bounded second moments of agent-generated pseudo-gradients, (ii) smoothness (or Lipschitz continuity) of the empirical loss function in the semantic embedding space, (iii) unbiased sampling induced by the bandit-based candidate selection mechanism.
+Under these assumptions, MAPGD admits an interpretation as a stochastic approximation process and achieves a sublinear convergence rate of $\mathcal{O}(1/\sqrt{T})$ in both convex and non-convex settings.
+Intuitively, the core components of MAPGD play complementary roles in ensuring stable optimization.
+HCGC aggregates semantically aligned refinement directions and mitigates conflicting updates, reducing variance in the fused pseudo-gradient.
+CAAW further emphasizes agents that have demonstrated consistent utility across iterations.
+Finally, bandit-based candidate selection maintains sufficient exploration while preventing systematic bias.
+Together, these mechanisms ensure that the overall optimization dynamics remain stable and progressively improve prompt quality over iterations.
+4 Experiments
+Experiments evaluate MAPGD in terms of effectiveness, efficiency, and robustness.
+For fair comparison, we reproduce ProTeGi under identical settings and further conduct ablations and token consumption analysis.
+A real-world text-generation case study is also included (Appendix I) to demonstrate applicability beyond benchmarks.
+4.1 Experimental Details
+Datasets.
+Following ProTeGi’s protocol, we evaluate on four classification benchmarks (Jailbreak, Ethos, LIAR, Sarcasm) and three arithmetic reasoning datasets (GSM8K, AQUARAT, SVAMP).
+For each dataset, we sample 50/150 instances for dev/test and report test F1 (classification) or accuracy (reasoning).
+Dataset details are in Appendix A.
+Setup.
+We use the January 2023 gpt-3.5-turbo via Azure OpenAI unless stated otherwise (temperature $0.0$ for classification).
+Results are averaged over three runs; F1 is computed by max-pooling over the final prompt beam.
+All methods share identical initial prompts, data, and random seeds, with a fixed budget of T=10T{=}10 and default hyperparameters.
+Unless stated otherwise, MAPGD follows Section 3.3 with $\phi{=}\texttt{all-MiniLM-L6-v2}$, $\theta_{\text{conflict}}{=}0.3$, clustering threshold $0.7$, $\theta_{\text{diversity}}{=}0.7$, and UCB selection with 80 evaluations; $\lambda{=}1$ (Appendix D).
+All experiments in Section 4.2 use these modules unless ablated.
+Baselines.
+We primarily compare against ProTeGi Pryzant et al.
+(2023) under identical query budgets and initialization, and additionally include MC search, RL-style edit methods, AutoGPT, and three reasoning baselines (InsZero, Instinct, PromptWizard).
+Full descriptions are in Appendix B.
+4.2 Experimental Results
+Figure 3: Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks (evaluated up to 50 queries for fair comparison with ProTeGi).
+Classification benchmarks.
+Figure 3 reports test F1 under different per-candidate evaluation budgets (up to 50 queries, following ProTeGi for fairness).
+MAPGD consistently achieves the best performance across all four datasets, with most gains emerging within the first 30–40 evaluations.
+The improvements are most pronounced on conflict-prone tasks, suggesting that explicit conflict disentanglement and adaptive coordination are critical when refinement signals disagree: on Jailbreak, MAPGD reaches 0.880 at 50 queries (vs.
+0.833 for ProTeGi), and on LIAR it attains 0.710 (vs.
+0.660).
+We attribute these gains to HCGC, which clusters compatible pseudo-gradients before fusion to reduce semantic interference, and CAAW, which down-weights unreliable agents to prevent noisy edits from dominating early iterations, thereby improving budget efficiency.
+On the more saturated Ethos dataset, MAPGD still improves to 0.980 (vs.
+0.965), indicating that coordinated multi-agent feedback can extract residual refinements even near a performance ceiling.
+We follow the 50-query budget for strict comparability; extended-budget results are deferred to Appendix C.
+Arithmetic reasoning benchmarks.
+Table 1 evaluates MAPGD on GSM8k, AQUARAT, and SVAMP, compared against InsZero, Instinct, and PromptWizard.
+MAPGD achieves consistent gains across all three datasets, most notably on GSM8k (93.5%), improving over the strongest baseline by +3.5 points.
+We hypothesize that the benefit on reasoning stems from the same coordination principle: specialized agents propose complementary edits (e.g., step-by-step constraints, format strictness, and exemplar guidance), while HCGC/CAAW suppress incompatible or low-yield directions, leading to more stable chain-of-thought prompting.
+MAPGD also improves the best baseline on AQUARAT (60.3% vs.
+58.2%) and SVAMP (84.1% vs.
+82.3%), confirming transfer beyond classification.
+Method GSM8k AQUARAT SVAMP
+InsZero 74.2 54.3 79.5
+Instinct 74.5 54.7 81.0
+PromptWizard 90.0 58.2 82.3
+[…]
+Table 1: Arithmetic reasoning accuracy (%) on GSM8k, AQUARAT, and SVAMP. Best results in bold.
+4.3 Ablation Study
+To isolate the sources of MAPGD’s gains, we conduct a component-wise ablation by selectively disabling HCGC and CAAW.
+Table 2 reports four settings (I–IV), where Setting I removes both modules and serves as a minimal baseline.
+Enabling either module leads to consistent improvements.
+Introducing HCGC (III vs.
+I) yields systematic gains, confirming that geometry-aware clustering prior to fusion is essential for reducing semantic interference among pseudo-gradients.
+CAAW alone (II vs.
+I) also improves performance, indicating that uniform averaging is suboptimal under varying agent reliability; adaptive reweighting based on validation feedback enhances robustness.
+The combined setting (IV) consistently outperforms either component alone, suggesting strong complementarity: HCGC resolves directional incompatibility via conflict disentanglement, whereas CAAW addresses reliability mismatch through adaptive credit assignment.
+Gains are smaller on saturated datasets (e.g., Ethos), consistent with ceiling effects, while conflict-prone tasks benefit more from structured coordination.
+Table 2: Ablation study on core MAPGD components. $\checkmark$ indicates the component is enabled. Results are F1 score (mean over 3 runs).
+Setting HCGC CAAW Jailbreak Ethos LIAR Sarcasm
+I $\times$ $\times$ 0.82 0.91 0.62 0.83
+II $\times$ $\checkmark$ 0.84 0.94 0.62 0.87
+III $\checkmark$ $\times$ 0.85 0.95 0.63 0.88
+IV $\checkmark$ $\checkmark$ 0.88 0.98 0.65 0.91
+4.4 Robustness to Dataset Sampling Size
+Following the standard ProTeGi evaluation protocol, our main results use 50 development and 150 test examples per dataset to ensure direct comparability.
+To verify that MAPGD’s gains are not an artifact of small evaluation sets, we additionally evaluate larger subsamples on three representative classification benchmarks.
+Table 3 reports results under three sampling regimes (50/150, 100/300, and 200/500).
+MAPGD consistently outperforms ProTeGi and MC across all datasets and sampling sizes, indicating that its improvements persist as the evaluation set grows.
+Dataset Method 50/150 100/300 200/500
+Jailbreak ProTeGi 0.82 0.83 0.81
+MC 0.76 0.77 0.75
+[…]
+Ethos ProTeGi 0.93 0.94 0.94
+MC 0.89 0.90 0.90
+[…]
+LIAR ProTeGi 0.65 0.62 0.63
+MC 0.62 0.63 0.63
+[…]
+Table 3: Robustness to dataset sampling size (F1 score) under three dev/test splits.
+4.5 Token Consumption Analysis
+Token efficiency is evaluated under the setup of Section 4.1 by counting both input and output tokens per API call.
+We report total tokens, number of calls, and performance-per-token.
+As shown in Table 4, MAPGD attains higher F1 with fewer tokens and calls than ProTeGi, reducing per-iteration token cost by $\sim$8% and improving performance-per-token by >>10%, suggesting reduced redundant queries via HCGC and CAAW.
+Metric ProTeGi MAPGD (full)
+[…]
+[…]
+[…]
+[…]
+Table 4: Token consumption comparison between ProTeGi and MAPGD (averaged over 3 runs). MAPGD achieves lower token usage and higher efficiency.
+4.6 Computational Overhead Analysis
+We compare runtime overhead against ProTeGi under identical settings.
+MAPGD shares evaluation, prompt expansion, and bandit selection with ProTeGi, but adds multi-agent gradient generation and embedding-based clustering.
+As shown in Table 5, the extra overhead mainly comes from multi-agent gradient generation, which scales linearly with the number of agents.
+Embedding and HCGC incur an almost constant cost, and with four agents the overall overhead remains a small fraction of total runtime.
+Component ProTeGi (s) MAPGD (s)
+Gradient generation 10.2 37.9
+Embedding + HCGC 24.4
+Prompt expansion 33.7 28.7
+Bandit selection 30.3 26.5
+Other 85.2 84.0
+[…]
+Table 5: Runtime breakdown (in seconds) comparing ProTeGi and MAPGD.
+4.7 Agent Number Sensitivity
+To motivate the use of four specialized agents, we vary the number of agents $N\in\{2,4,6\}$ under the setup of Section 4.1.
+For $N=2$, only the Instruction Specialist and Example Curator are used; for $N=6$, two additional style/format agents are added.
+As shown in Figure 4, average F1 improves substantially from $N=2$ to $N=4$, while $N=6$ yields only marginal gains with higher token cost, indicating diminishing returns.
+Thus, $N=4$ offers the best effectiveness–efficiency trade-off.
+Figure 4: Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
+4.8 Experimental Insights
+MAPGD consistently outperforms single-agent ProTeGi and other baselines, with the largest gains on LIAR and Jailbreak where refinement signals often conflict.
+Ablations show HCGC and CAAW are complementary: removing either hurts performance, and removing both approaches the single-agent regime.
+MAPGD is also more token-efficient than ProTeGi (fewer API calls, >>10% higher performance-per-token) and transfers well to arithmetic reasoning, surpassing strong baselines on GSM8k, AQUARAT, and SVAMP.
+Appendix H provides a qualitative case study illustrating how HCGC and CAAW resolve conflicting pseudo-gradients.
+5 Conclusion
+This work introduces MAPGD, a coordinated multi-agent framework for prompt optimization that integrates HCGC and CAAW to resolve conflicting refinement directions and reduce the impact of unreliable agents.
+Across classification and arithmetic reasoning benchmarks, MAPGD consistently improves over strong baselines while lowering token consumption, and it admits a sublinear convergence guarantee of $O(1/\sqrt{T})$ under standard stochastic approximation assumptions (Appendix G).
+Overall, MAPGD provides an interpretable, robust, and resource-efficient approach to prompt optimization via structured multi-agent collaboration.
+Limitations
+MAPGD relies on semantic embeddings to detect conflicts and cluster pseudo-gradients; under domain shift or highly specialized jargon, similarity estimates may become unreliable and degrade coordination quality.
+The method also depends on reasonably informative agent feedback: if multiple agents are systematically misaligned, adaptive weighting may not fully suppress harmful edits.
+Finally, compared to single-agent methods, MAPGD introduces extra overhead from multi-agent generation and coordination (embedding/clustering), which may limit deployment under strict latency or budget constraints.
+Appendix A Datasets Details
 To ensure comparability with ProTeGi, we evaluate MAPGD on the same four benchmark NLP classification tasks that span diverse domains and languages:
@@
 In addition, we include a real-world text generation case study in the Appendix to highlight MAPGD’s applicability beyond controlled benchmarks.
-4.2. Experimental Setup
-For each dataset, we randomly sample 50 examples for development and 150 for testing.
-All reported results are averaged over three independent runs to mitigate variance.
-The evaluation metric is binary F1 score on the test set, computed by max-pooling over the final beam of candidate prompts.
-Unless otherwise stated, all experiments are conducted with the January 2023 version of gpt-3.5-turbo through the Azure OpenAI API.
-The temperature is set to 0.0 for classification tasks to ensure deterministic predictions.
-To ensure a fair comparison, all methods—including baselines and MAPGD—use the same initial prompts, training data, and random seeds.
-The iteration budget is fixed at $T=10$ across all experiments.
-We adopt default parameter values without additional hyperparameter search, to ensure that performance differences are attributable to algorithmic design rather than parameter tuning.
-For our MAPGD implementation, we instantiate the framework with specific hyperparameters introduced in Section 3.3.
-We use the all-MiniLM-L6-v2 Sentence Transformer as the gradient encoder $\phi$.
-The conflict threshold $\theta_{conflict}$ is set to 0.3, and we use a cluster similarity threshold of 0.7 for grouping gradients.
-The diversity threshold $\theta_{diversity}$ is set to 0.7 to ensure varied candidate prompts.
-Bandit-based selection adopts a UCB strategy with a budget of 80 evaluations.We set the CAAW weighting parameter $\lambda$ to 1 (see Appendix A for a sensitivity analysis on different values).
-Accordingly, all experiments reported in Section 4.4 were conducted with these modules enabled, unless explicitly ablated.
-4.3. Baselines
+Appendix B Baselines Details
 To evaluate the effectiveness of MAPGD, we compare it against a set of non-parametric prompt optimization methods, following the setup of ProTeGi Pryzant et al.
@@
 This collection of baselines allows us to evaluate MAPGD not only against single-agent gradient-based optimization, but also against strong arithmetic reasoning methods optimized for chain-of-thought style prompting.
-4.4. Experimental Results
-Figure 5 summarizes the overall performance across the four benchmark classification datasets.
-Our results demonstrate that MAPGD consistently achieves the best performance under varying evaluation budgets, outperforming all baselines by a notable margin.
-On the Jailbreak dataset, MAPGD attains an average F1 of 0.86, surpassing ProTeGi (0.82) and MC (0.76), while substantially improving over RL (0.67) and AutoGPT (0.62).
-The improvement margin remains stable across different budgets, reaching up to +5.5% over ProTeGi and +10.9% over MC at 30 evaluations.
-For Ethos, MAPGD pushes performance from the strong baseline of 0.93 ($p_{0}$) to 0.98, showing a consistent +1.5% advantage over ProTeGi and +4.0% over MC.
-These improvements highlight MAPGD’s ability to leverage gradient fusion even in relatively saturated tasks.
-On the more challenging LIAR dataset, MAPGD achieves 0.71, representing the largest relative gains: +8.0% over ProTeGi, +11.0% over MC, and a substantial +17.0% over the initial prompt.
-This suggests that multi-agent collaboration and semantic clustering are particularly beneficial in noisy, fact-checking tasks.
-For Sarcasm, MAPGD steadily improves from 0.84 ($p_{0}$) to 0.91, outperforming ProTeGi (0.87) and MC (0.85).
-Interestingly, AutoGPT sometimes reduces performance below the initial prompt, echoing earlier findings that unguided agent feedback may destabilize optimization.
-Beyond classification, we further evaluate MAPGD on arithmetic reasoning benchmarks, which require detailed multi-step inference.
-Table 1 compares MAPGD against three strong baselines—InsZero, Instinct, and PromptWizard—on GSM8k, AQUARAT, and SVAMP.
-MAPGD achieves consistent improvements across all datasets, with particularly notable gains on GSM8k (93.5%), surpassing the previous best by +3.5 points.
-On AQUARAT and SVAMP, MAPGD reaches 60.3% and 84.1%, respectively, representing stable improvements over the strongest baselines.
-These results confirm that MAPGD not only excels in classification but also transfers effectively to challenging reasoning tasks.
-Overall, MAPGD improves over ProTeGi and MC by 3–7% on average for classification, while maintaining strong gains on arithmetic reasoning datasets.
-These results validate our hypothesis that multi-agent specialization and hypersphere-constrained clustering provide more robust and adaptive optimization compared to single-agent or rule-based approaches.
-Table 1. Arithmetic reasoning accuracy (%) on GSM8k, AQUARAT, and SVAMP. Best results in bold.
-Method GSM8k AQUARAT SVAMP
-InsZero 74.2 54.3 79.5
-Instinct 74.5 54.7 81.0
-PromptWizard 90.0 58.2 82.3
-MAPGD 93.5 60.3 84.1
-Figure 5. Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks.
-[…]
-Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks.
-4.5. Ablation Study
-To precisely attribute the performance gains of MAPGD, we perform a comprehensive ablation study (see Table 2 for the detailed configurations and results).
-Beyond standard ablation variants (e.g., No Gradient Fusion, Single-Agent Only, No Specialization), we design targeted experiments specifically to isolate the contributions of the two proposed modules HCGC and CAAW, as well as to evaluate resource efficiency and robustness.
-Ablation configurations.
-We evaluate the following variants:
-  1. (1)
-    MAPGD (full): our full method with HCGC and CAAW.
-  2. (2)
-    No Gradient Fusion: disable the fusion stage; agents act independently and their suggestions are applied without semantic clustering.
-  3. (3)
-    Single-Agent Only: only one specialized agent (instruction specialist) is used.
-  4. (4)
-    No Specialization: agent roles are randomized each iteration (no fixed expertise).
-  5. (5)
-    w/o HCGC: keep clustering but remove the hypersphere angular-margin constraint.
-This reduces to standard K-means clustering with cosine similarity, which acts as an off-the-shelf baseline for comparison.
-  6. (6)
-    w/o CAAW: replace Channel-Adaptive Agent Weighting with simple uniform averaging of agent gradients.
-  7. (7)
-    w/o HCGC+CAAW: use standard K-means clustering with cosine similarity (no angular margin) and replace adaptive weighting with uniform fusion.
-Table 2. Ablation study: classification performance (F1 score, mean $\pm$ std over 3 runs). Best results in bold.
-Method Jailbreak Ethos LIAR Sarcasm
-MAPGD (full) 0.88 $\pm$ 0.01 0.98 $\pm$ 0.02 0.71 $\pm$ 0.01 0.91 $\pm$ 0.01
-No Gradient Fusion 0.83 $\pm$ 0.02 0.92 $\pm$ 0.02 0.67 $\pm$ 0.02 0.86 $\pm$ 0.02
-Single Agent Only 0.80 $\pm$ 0.03 0.87 $\pm$ 0.03 0.62 $\pm$ 0.02 0.82 $\pm$ 0.03
-No Specialization 0.80 $\pm$ 0.02 0.90 $\pm$ 0.02 0.65 $\pm$ 0.02 0.84 $\pm$ 0.02
-w/o HCGC 0.84 $\pm$ 0.02 0.94 $\pm$ 0.02 0.68 $\pm$ 0.02 0.87 $\pm$ 0.02
-w/o CAAW 0.85 $\pm$ 0.02 0.95 $\pm$ 0.02 0.69 $\pm$ 0.02 0.88 $\pm$ 0.02
-w/o HCGC + CAAW 0.82 $\pm$ 0.03 0.91 $\pm$ 0.03 0.66 $\pm$ 0.02 0.83 $\pm$ 0.02
-4.6. Token Consumption Analysis
-To further assess the efficiency of MAPGD, we conducted a dedicated token consumption experiment, following the same setup as in Section 4.2.
-In this experiment, we recorded the number of tokens used in every API call, where the token cost of a single call is defined as the sum of input tokens and output tokens.
-For each method, we accumulated token usage across all iterations and runs, and report the total tokens, average tokens per iteration, and total number of LLM calls.
-In addition, we report performance efficiency as the ratio of average F1 score to total token usage.
-Table 3 presents the results, comparing MAPGD with the ProTeGi.
-MAPGD not only achieves higher F1 scores, but also requires fewer tokens and API calls overall.
-Specifically, MAPGD reduces the average token cost per iteration from roughly 25.6k to 23.6k, corresponding to a $\sim$8% saving in token usage.
-At the same time, MAPGD improves the performance-per-token efficiency by more than 10%.
-These findings indicate that multi-agent collaboration with HCGC and CAAW enables more effective use of query budgets, reducing redundancy while preserving accuracy.
-Table 3. Token consumption comparison between ProTeGi and MAPGD (averaged over 3 runs). MAPGD achieves lower token usage and higher efficiency.
-Metric ProTeGi MAPGD (full)
-Total Tokens 256k 236k
-Calls 962 643
-Avg F1 0.83 0.87
-F1 / Token ($\times 10^{-6}$) 3.24 3.69
-4.7. Agent Number Sensitivity
-To better justify the design choice of using four specialized agents, we conduct a sensitivity analysis by varying the number of agents $N\in\{2,4,6\}$.
-For $N=2$, we retain only the Instruction Specialist and Example Curator; for $N=6$, we extend the framework with two additional stylistic and format-focused agents.
-All other settings follow Section 4.2.
-Figure 6 shows the average F1 score across the four datasets.
-We observe that performance improves substantially when moving from $N=2$ to $N=4$, confirming the benefit of multi-agent specialization.
-Increasing to $N=6$ yields only marginal gains while incurring higher token cost, suggesting diminishing returns.
-These results indicate that four agents provide a favorable balance between effectiveness and efficiency, validating our default configuration.
-Overall, these findings suggest that an excessively small agent set cannot comprehensively capture different prompt dimensions, whereas an overly large set introduces redundancy and potential conflicts.
-A moderate number of agents, four in our case, provides the best trade-off.
-Figure 6. Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
-[…]
-Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
-4.8. Experimental Insights
-Our experimental evaluation provides several key insights:
-  • Effectiveness of multi-agent collaboration.
-Across all four benchmark tasks, MAPGD consistently outperforms single-agent ProTeGi and other baselines.
-The largest gains are observed on LIAR and Jailbreak, where conflicting gradient signals are common.
-This demonstrates that multi-agent specialization, combined with semantic clustering, can effectively resolve gradient conflicts and guide prompt optimization toward more robust improvements.
-  • Contributions of HCGC and CAAW.
-The ablation study confirms that both proposed modules are indispensable.
-Removing HCGC reduces the ability to enforce semantic compactness and separation, leading to less stable optimization.
-Disabling CAAW degrades robustness by failing to downweight weaker agents.
-When both modules are removed, MAPGD collapses close to single-agent performance, highlighting their complementary roles.
-  • Efficiency in token usage.
-Despite employing multiple agents, MAPGD achieves lower overall token consumption and fewer API calls than ProTeGi.
-This efficiency stems from clustering redundant gradients and adaptively weighting agents, which reduce unnecessary evaluations.
-As a result, MAPGD achieves over 10% higher performance-per-token efficiency, indicating that it not only improves accuracy but also makes more economical use of query budgets.
-  • Robustness across tasks.
-The improvements are consistent across datasets of varying size, domain, and language, suggesting that MAPGD generalizes well to heterogeneous NLP tasks.
-Furthermore, the case study in the Appendix illustrates its applicability to open-ended generation, extending beyond controlled classification benchmarks.
-  • Transferability to reasoning tasks.
-MAPGD also demonstrates strong performance on arithmetic reasoning datasets, surpassing competitive baselines such as PromptWizard.
-Its consistent improvements on GSM8k, AQUARAT, and SVAMP highlight that the proposed gradient-inspired multi-agent optimization is not limited to classification, but can effectively enhance chain-of-thought style reasoning as well.
-In summary, these insights show that MAPGD achieves a favorable balance between performance and efficiency, while its modular design (HCGC + CAAW) provides both effectiveness and robustness in diverse prompt optimization scenarios.
-5. Discussion
-MAPGD advances prompt optimization by combining gradient-inspired reasoning with multi-agent collaboration.
-Its two key mechanisms—Hypersphere-Constrained Gradient Clustering and Channel-Adaptive Agent Weighting enable more structured, interpretable, and efficient optimization.
-Modularity.
-By separating gradient generation, clustering, weighting, and candidate selection, MAPGD clarifies which components drive performance, supporting transparent analysis and principled extensions.
-Conflict resolution.
-HCGC enforces angular margin constraints on the hypersphere, promoting intra-cluster compactness and inter-cluster separation.
-This directly addresses conflicts among agent-generated gradients, common in tasks such as fake news detection or jailbreak classification, and parallels gradient surgery in multi-task learning but for discrete prompts.
-Adaptive collaboration.
-CAAW allocates influence based on agent reliability, reducing noise from weaker gradients while amplifying consistent signals.
-This is useful when tasks emphasize different reasoning dimensions (e.g., semantic consistency vs.
-contextual grounding).
-Budget-awareness. With lightweight bandit-based selection, MAPGD improves performance-per-token efficiency, achieving higher accuracy with fewer redundant API calls.
-Limitations.
-MAPGD depends on embedding quality and agent reliability; under strong domain shifts, clustering may be unstable and weighting may amplify spurious signals.
-Future work may strengthen semantic representations and improve agent calibration.
-Future directions. Promising avenues include:
-  1. (1)
-    Cross-task generalization: training reusable agents with transferable clustering and weighting behaviors.
-  2. (2)
-    Human-in-the-loop fusion: leveraging expert oversight to guide clustering and validate weighting.
-  3. (3)
-    Hybrid optimization: integrating MAPGD with differentiable prompt tuning to balance interpretability and efficiency.
-6. Conclusion
-We presented MAPGD: Multi-Agent Prompt Gradient Descent, a novel framework for optimizing prompts in large language models.
-The key innovations of MAPGD are HCGC, which resolves conflicts through hypersphere-constrained clustering, and CAAW, which adaptively balances the contributions of specialized agents.
-Together, these mechanisms enable diverse agents to collaborate effectively while maintaining semantic alignment.
-Empirically, MAPGD achieves consistent improvements over strong baselines such as ProTeGi, Monte Carlo search, reinforcement learning–based methods, and recent arithmetic reasoning frameworks.
-On four benchmark classification tasks, MAPGD improves F1 scores by 3–7% on average while reducing token consumption by $\sim$8%, leading to better performance-per-token efficiency.
-Beyond classification, MAPGD also demonstrates strong transferability to arithmetic reasoning datasets such as GSM8k, AQUARAT, and SVAMP, where it surpasses competitive baselines including PromptWizard.
-These results confirm that gradient-inspired multi-agent collaboration is broadly applicable across both discriminative and reasoning tasks.
-Theoretically, we show that MAPGD achieves a sublinear convergence rate of $O(1/\sqrt{T})$ under standard stochastic approximation assumptions (see Appendix C for details).
-This result indicates that even in a discrete prompt space, the semantic gradient mechanism—stabilized by HCGC and CAAW—retains the efficiency of classical stochastic gradient descent.
-While challenges remain, particularly regarding embedding robustness and LLM feedback reliability, MAPGD provides both a practical toolkit for efficient prompt optimization and a theoretical foundation linking discrete semantic updates with continuous optimization theory.
-Future work may extend HCGC and CAAW to multimodal prompts, incorporate human preference alignment, and develop cross-domain reusable agents.
-In summary, MAPGD demonstrates that multi-agent collaboration, when combined with principled clustering and adaptive weighting, leads to more robust, interpretable, and resource-efficient prompt optimization, with demonstrated effectiveness across classification, reasoning, and beyond.
-Appendix A CAAW Parameter Sensitivity
-The Channel-Adaptive Agent Weighting module uses a temperature parameter $\lambda$ to control the emphasis on historically effective agents during gradient fusion.
+Appendix C Extended Evaluation Budgets Beyond 50 Queries
+To verify whether our conclusions depend on the 50-query budget, we further extend the evaluation budget to 60 and 70 queries on three representative datasets (Jailbreak, LIAR, and Sarcasm).
+We keep the same evaluation protocol and report results in Table 6.
+Overall, the method ranking remains unchanged, and the additional gains beyond 50 are marginal, suggesting that a 50-query budget is sufficient to capture MAPGD’s performance while preserving fairness with prior work (e.g., ProTeGi).
+Dataset Budget ProTeGi MC MAPGD
+Jailbreak 50 0.833 0.773 0.880
+60 0.836 0.773 0.885
+70 0.842 0.773 0.887
+LIAR 50 0.660 0.630 0.710
+60 0.672 0.630 0.720
+70 0.681 0.630 0.725
+Sarcasm 50 0.876 0.853 0.913
+60 0.879 0.853 0.918
+70 0.880 0.853 0.920
+Table 6: Performance under extended evaluation budgets (60/70 queries). Gains beyond 50 queries are small and do not change the ranking among methods.
+Across these datasets, MAPGD improves by only +0.007 (Jailbreak), +0.015 (LIAR), and +0.007 (Sarcasm) from 50 to 70 queries, confirming a clear diminishing-return trend after approximately 30–40 evaluations, while leaving the main conclusions unchanged.
+Appendix D CAAW Parameter Sensitivity
+The CAAW module uses a temperature parameter $\lambda$ to control the emphasis on historically effective agents during gradient fusion.
 To evaluate the robustness of MAPGD with respect to $\lambda$, we conducted experiments on two representative datasets: LIAR(binary) and GSM8K (mathematical reasoning).
@@
 The results indicate that moderate variations of $\lambda$ do not significantly affect the overall performance, supporting the choice of $\lambda=1$ in the main experiments.
-Table 4. Test accuracy of MAPGD under different CAAW weighting parameter $\lambda$. Performance differences are minor, demonstrating robustness to $\lambda$.
+Table 7: Test accuracy of MAPGD under different CAAW weighting parameter $\lambda$. Performance differences are minor, demonstrating robustness to $\lambda$.
 $\lambda$ LIAR Accuracy GSM8K Accuracy
@@
 These variations are small relative to the overall performance gains achieved by multi-agent gradient fusion, indicating that while $\lambda$ influences the sharpness of agent weighting, MAPGD’s performance remains stable for reasonable choices around the default setting of $\lambda=1$.
-Appendix B Complexity and Parallelism
+Appendix E Comparison with Recent Concurrent Work
+We are aware of two recent concurrent works that are closely related in spirit to prompt optimization with multi-agent components, namely MAPRO Zhang et al.
+(2025) and MA-SAPO Seo et al.
+(2025).
+While these works share a high-level motivation with MAPGD, we did not include them as baselines in our main experimental comparisons due to comparability constraints in problem setting and resource assumptions.
+MAPRO Zhang et al. (2025).
+MAPRO studies prompt optimization within an explicit multi-agent system (MAS): it jointly configures role prompts for multiple agents and leverages the MAS interaction topology (e.g., DAG-structured communication) together with node-/edge-level reward modeling and global inference to select prompt combinations.
+In contrast, MAPGD targets task-level prompt optimization under a fixed query budget, where specialized agents propose pseudo-gradients that are subsequently coordinated (via HCGC and CAAW) to iteratively refine a single task prompt.
+Because MAPRO’s formulation depends on an explicit MAS graph, per-agent prompts, and interaction-specific rewards, directly porting it to our benchmark protocol would require redesigning the environment and supervision signals, making a head-to-head comparison not directly aligned with our evaluation setting.
+MA-SAPO Seo et al. (2025).
+MA-SAPO optimizes prompts through retrieved reasoning assets: it constructs a searchable library of semi-structured “reasoning cards” from labeled or scored prompt–response data, and then performs analysis–refinement guided by top-$k$ retrieved assets at test time.
+This introduces an additional external memory/resource assumption (i.e., an offline asset corpus and its construction pipeline) that is orthogonal to our training-free, black-box optimization protocol.
+In our evaluation, we intentionally restrict methods to operate under the same query budget without relying on extra annotated corpora or a separately curated retrieval memory.
+Under these constraints, implementing MA-SAPO would either (i) require additional data collection/annotation to build assets, or (ii) reuse task-specific optimization traces as assets, which changes the method definition and complicates fairness.
+Summary and future work.
+Overall, MAPRO and MA-SAPO represent complementary directions—MAS-level joint role-prompt configuration and retrieval-augmented prompt rewriting—whereas MAPGD focuses on multi-agent pseudo-gradient coordination for budgeted task-level prompt refinement.
+Integrating these paradigms into a unified and strictly comparable evaluation (e.g., extending MAPGD with optional retrieval memories or evaluating MAPRO-style joint prompting on explicit MAS benchmarks) is an important direction for future work.
+Appendix F Complexity and Parallelism
 The computational bottleneck of MAPGD with HCGC and CAAW (see Algorithm 1) lies in LLM calls for gradient generation, cluster fusion, and prompt expansion, whereas embedding, clustering, similarity checks, and adaptive weighting involve comparatively lightweight vector operations.
@@
 Stage costs (time).
-  1. (1)
+  1. 1.
     Agent gradient generation: Each of the $N$ agents generates $m$ gradients using LLMs, giving $O(N)$ LLM calls. Parallelization reduces wall time to $\approx\max t_{\text{LLM}}$.
-  2. (2)
+  2. 2.
     HCGC embedding + conflict detection: Embedding each gradient costs $O(Gd)$; computing pairwise angular conflicts costs $O(G^{2}d)$.
 For small $G$ (e.g., $G\sim 16$), this is negligible compared with LLM calls.
-  3. (3)
+  3. 3.
     HCGC clustering: KMeans clustering over $G$ gradients into $K$ clusters with $I$ iterations costs $O(GKId)$, again minor due to small $G$.
-  4. (4)
+  4. 4.
     HCGC fusion (LLM): Clusters with multiple gradients require one LLM call per cluster to fuse into a coherent gradient. Total LLM calls up to $O(K_{\text{merge}})$, parallelizable across clusters.
-  5. (5)
+  5. 5.
     CAAW weighting and fusion: Per-cluster adaptive weighting is $O(G)$ vector operations, negligible compared to LLM fusion.
-  6. (6)
+  6. 6.
     Prompt expansion + MC paraphrasing: Applying each fused $O(|\tilde{\mathcal{G}}|) gradient$ applications. MC paraphrasing for each variant: $O(|\tilde{\mathcal{G}}|sn_{\text{mc}})$ LLM calls.
-  7. (7)
+  7. 7.
     Diversity filtering: Computing embeddings: $O(|\mathcal{C}|d)$; naive pairwise similarities: $O(|\mathcal{C}|^{2})$. For tens of candidate prompts, this remains negligible.
-  8. (8)
+  8. 8.
     Bandit evaluation: Probing $K_{\text{eval}}$ arms over mini-batches of size $b$ for $T_{b}$ rounds: $O(K_{\text{eval}}bT_{b})$, significantly cheaper than exhaustive evaluation $O(|\mathcal{C}|\,|D_{\text{dev}}|)$.
@@
 Optional caches (unique prompts or intermediate LLM outputs) scale linearly with the number of prompts and clusters.
-Algorithm 1 MAPGD with HCGC and CAAW
-1:Initial prompt $p_{0}$, train data $D_{\text{train}}$, dev data $D_{\text{dev}}$, agents $\{A_{i}\}_{i=1}^{N}$, iterations $R$, beam width $k$
-2:Initialize best prompt $p_{\star}^{(0)}\leftarrow p_{0}$, beam $B_{0}\leftarrow\{p_{0}\}$; initialize each agent prompt $A_{i}.p\leftarrow p_{0}$
-3:for $t=1$ to $R$ do
-4:  $M_{t}\leftarrow\textsc{SampleMiniBatch}(D_{\text{train}},b)$
-5:  $\mathcal{G}_{t}\leftarrow\textsc{GenerateAgentGradients}(\{A_{i}\},M_{t})$ $\triangleright$ Alg. 2
-6:  $\tilde{\mathcal{G}}_{t}\leftarrow\textsc{HCGC}(\mathcal{G}_{t})$ $\triangleright$ Alg. 3
-7:  $F_{t}\leftarrow\textsc{CAAWFuse}(\tilde{\mathcal{G}}_{t},D_{\text{dev}})$ $\triangleright$ Alg. 4
-8:  $\mathcal{C}_{t}\leftarrow\textsc{ExpandPrompts}(p_{\star}^{(t-1)},F_{t})$
-9:  $B_{t},p_{\star}^{(t)}\leftarrow\textsc{BanditSelect}(\mathcal{C}_{t},D_{\text{dev}},k)$
-[…]
-[…]
-12:  end if
-13:end for
-14:return Best prompt over $\{p_{\star}^{(1)},\dots,p_{\star}^{(t)}\}$
 Algorithm 2 Multi-Agent Textual Gradient Generation
@@
 7:return Updated agents $\{A_{i}\}$
-Appendix C Theoretical Analysis
+Appendix G Theoretical Analysis
 In this section, we provide convergence guarantees for MAPGD under mild assumptions, following the stochastic approximation framework.
@@
 We show that, despite operating in a structured and discrete search space, MAPGD achieves the same sublinear convergence rate of $O(1/\sqrt{T})$ in both convex and non-convex settings.
-C.1. Assumptions
+G.1 Assumptions
 We begin with a set of assumptions standard in stochastic optimization but reinterpreted in the context of multi-agent prompt optimization.
@@
   • (A3) Smoothness or Lipschitzness. For convex tasks, $F$ is $G$-Lipschitz with domain diameter $D$. For non-convex tasks, $F$ is $L$-smooth: $\|\nabla F(u)-\nabla F(v)\|\leq L\|u-v\|$.
-C.2. Supporting Lemmas
+G.2 Supporting Lemmas
 We restate two standard lemmas, adapted to the MAPGD setting.
@@
 satisfies
-$\|p^{(t+1)}-p^{*}\|^{2}\;\leq\;\|p^{(t)}-p^{*}\|^{2}-2\eta\langle g^{(t)},p^{(t)}-p^{*}\rangle+\eta^{2}\|g^{(t)}\|^{2}.$
+$\|p^{(t+1)}-p^{*}\|^{2}$ $\leq\|p^{(t)}-p^{*}\|^{2}-2\eta\langle g^{(t)},\,p^{(t)}-p^{*}\rangle$
+$\quad+\eta^{2}\|g^{(t)}\|^{2}.$
 Lemma 2 (Non-Convex Descent Lemma).
@@
 $F(p^{(t+1)})\;\leq\;F(p^{(t)})-\eta\langle\nabla F(p^{(t)}),g^{(t)}\rangle+\frac{L}{2}\eta^{2}\|g^{(t)}\|^{2}.$
-C.3. Main Results
+G.3 Main Results
 Convex Convergence.
@@
 Summing over $t=1\ldots T$ gives
-$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left[\|\nabla F(p^{(t)})\|^{2}\right]\;\leq\;\frac{2(F(p^{(1)})-F_{\inf})}{\mu T\eta}+\frac{L\sigma^{2}}{\mu}\eta.$
+$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left[\|\nabla F(p^{(t)})\|^{2}\right]\;\leq$ $\frac{2\big(F(p^{(1)})-F_{\inf}\big)}{\mu T\eta}$
+$\;+\;\frac{L\sigma^{2}}{\mu}\eta.$
 Balancing terms with $\eta=\Theta(1/\sqrt{T})$ yields the claimed rate.
-C.4. Summary of Theoretical Guarantees
+G.4 Summary of Theoretical Guarantees
 In this work, we bridge the gap between classical stochastic optimization, which assumes continuous parameter spaces, and the inherently discrete, structured nature of prompt optimization.
@@
 The argument unfolds along three main dimensions:
-  • Directional Alignment (A1): Although textual pseudo-gradients are not true mathematical gradients, Hypersphere-Constrained Gradient Clustering aggregates semantically coherent suggestions from multiple agents into a fused gradient that statistically aligns with the true descent direction.
+  • Directional Alignment (A1): Although textual pseudo-gradients are not true mathematical gradients, HCGC aggregates semantically coherent suggestions from multiple agents into a fused gradient that statistically aligns with the true descent direction.
   • Variance Control (A2): Discrete prompt edits are prone to high variance, but the bandit-based candidate selection mechanism filters out unreliable or detrimental candidates.
@@
 Consequently, despite operating in a discrete textual space, MAPGD achieves a robust sublinear convergence rate of $O(1/\sqrt{T})$, providing a solid theoretical foundation for the observed empirical effectiveness of our multi-agent, geometry-aware approach to prompt optimization.
-Appendix D Case Study: System Prompt Optimization
+Appendix H Qualitative Case Study: From Conflicting Pseudo-Gradients to Prompt Updates
+H.1 Case Setup
+We select a representative optimization step on the LIAR task where multiple specialized agents propose semantically conflicting pseudo-gradients.
+This step exhibits clear disagreement among agents while still leading to a measurable improvement in downstream validation behavior after gradient fusion.
+Item Value
+Task LIAR
+Iteration $t=3$
+Agents 4
+Raw Gradients 16
+Illustrated Conflict Pairs 2
+HCGC Clusters 3
+At this iteration, agents disagree on whether errors primarily stem from ambiguous label semantics, missing contextual cues, or insufficient decision boundary clarification, making it a suitable case for qualitative inspection.
+H.2 Raw Pseudo-Gradients from Specialized Agents
+We focus on two representative conflict pairs extracted from the same optimization step, which illustrate structurally different failure diagnoses and motivate the need for explicit conflict handling.
+Conflict Pair 1: Format Clarity vs Example Diversity.
+Format Designer:
+<START> The prompt instructs the model to answer with ‘Yes’ or ‘No’, but it does not specify what these labels represent.
+This ambiguity can cause the model to misinterpret the classification objective.
+<END>
+Example Curator:
+<START> The provided examples do not cover a sufficiently diverse range of statements. Including more varied and representative cases could help the model generalize better. <END>
+The cosine similarity between these two gradients is $0.13$ (angle =1.44=1.44 rad), indicating strong semantic disagreement.
+Importantly, the conflict is not superficial: the format-oriented gradient advocates tightening the input–output contract through explicit label semantics, while the example-oriented gradient pushes the prompt toward structural expansion via additional demonstrations.
+Applying both updates independently would lead to incompatible prompt modifications—one emphasizing deterministic label interpretation and the other reallocating prompt capacity to example coverage.
+Conflict Pair 2: Example Coverage vs Definition of “Lie”.
+Example Curator:
+<START> The examples fail to illustrate borderline or ambiguous cases, which may limit the model’s ability to generalize across different claim types. <END>
+Instruction Specialist:
+<START> The prompt does not clearly define what constitutes a ‘lie’ versus an inaccurate or exaggerated statement, leading to inconsistent decisions. <END>
+This pair exhibits a cosine similarity of $0.23$ (angle =1.34=1.34 rad), reflecting a fundamental disagreement about the source of errors.
+The example-focused gradient assumes that misclassification primarily stems from insufficient empirical coverage, whereas the instruction- focused gradient attributes failures to an underspecified decision criterion.
+These two perspectives induce different optimization directions: expanding illustrative evidence versus formalizing the semantic boundary of the task itself.
+Discussion.
+Both conflict pairs arise from valid but incompatible diagnostic assumptions.
+Crucially, neither gradient is inherently incorrect; each targets a distinct aspect of model behavior.
+This highlights the core challenge addressed by MAPGD: without explicit conflict detection and coordination, naïvely aggregating such gradients would obscure their semantic tension and lead to unstable or diluted prompt updates.
+H.3 Conflict Detection and HCGC Clustering
+Based on the two conflict pairs in §H.2, we observe clear semantic disagreement among agent-proposed directions.
+In particular, the Format Clarity vs Example Diversity pair has cosine similarity $0.13$ (angle =1.44=1.44 rad), and the Example Coverage vs Definition of “Lie” pair has cosine similarity $0.23$ (angle =1.34=1.34 rad).
+The corresponding angles ($1.44$ and $1.34$ rad) indicate substantial directional mismatch, and both similarities fall below the conflict threshold.
+HCGC therefore marks them as conflicting and avoids naïve averaging.
+After projecting all gradient embeddings onto the unit hypersphere and applying the angular-margin constraints, HCGC partitions the 16 raw pseudo-gradients into three coherent clusters, as summarized in Table 8.
+Table 8: HCGC clustering summary for the representative LIAR optimization step (t=3t{=}3).
+Cluster Agents Involved Semantic Theme
+C1 instruction, format label semantics and task definition (binding Yes/No to True/False, clarifying what is being judged)
+C2 example, style coverage and contextual cues for borderline cases (exaggeration, unverifiable framing, pragmatic signals)
+C3 instruction only decision-boundary calibration (strictness vs recall; discouraging overly permissive acceptance of absolute claims)
+This clustering reflects a key property of HCGC: conflicting gradients are not forced into a single averaged direction.
+Instead, they are separated into semantically consistent groups (Table 8), enabling subsequent fusion to preserve complementary information (e.g., clarifying label semantics in C1 while retaining coverage-oriented refinements in C2) without letting one channel dominate the other.
+The semantic themes are manually summarized by the authors to reflect the dominant intent of each cluster, rather than being directly emitted by the model.
+H.4 Channel-Adaptive Agent Weighting
+Given the clustered directions above, CAAW assigns each agent a contribution weight based on its recent performance signal, so that the subsequent fusion emphasizes gradients that are empirically reliable for the current task state.
+At this iteration, the adaptive weights are:
+[…]
+[…]
+These weights are consistent with the observed conflict structure.
+Clusters C1 and C2 correspond to two structurally competing but both valid refinement axes: (i) clarifying label semantics and the task definition (C1), and (ii) improving coverage and contextual handling of borderline cases (C2).
+CAAW prioritizes instruction- and example-driven gradients while down-weighting stylistic adjustments, preventing prompt capacity from being consumed by surface-level phrasing when the dominant failures are semantic (label meaning) and epistemic (what constitutes a lie).
+H.5 Fused Gradient and Prompt Modification
+After separating conflicting directions with HCGC and weighting them with CAAW, MAPGD fuses the cluster-level signals into a single coherent update that preserves complementary intent across clusters.
+In this case, the fused direction primarily integrates (a) C1: explicit label semantics and task definition, and (b) C2: guidance for borderline or unverifiable claims, while avoiding excessive prompt expansion.
+Before:
+“The following statement is either true or false. Answer Yes or No.”
+After:
+“Determine whether the statement is factually accurate.
+Treat Yes as True and No as False.
+If the statement relies on exaggeration, satire, or unverifiable claims, judge it as False.
+Return only Yes or No.”
+This modification reflects the fused gradient in two concrete ways: (i) it resolves the format/label ambiguity highlighted in C1 by explicitly binding Yes/No to True/False, and (ii) it incorporates the borderline-case handling emphasized by C2 without requiring lengthy demonstrations, thereby avoiding the prompt-structure conflict that would arise from naïvely mixing rule-heavy and example-heavy updates.
+H.6 Behavioral Change on Previously Misclassified Examples
+We report qualitative behavioral changes on examples that were previously misclassified due to (i) label ambiguity and (ii) overly permissive interpretation of absolute or exaggerated claims.
+Example 1 (Absolute claim / overgeneralization).
+Claim: “The senator has voted against every environmental bill.”
+Before: Yes   After: No
+The updated prompt instructs the model to treat unverifiable or sweeping claims as False, which directly corrects the previous tendency to accept absolute statements (e.g., “every”) as True.
+Example 2 (Borderline / unverifiable framing).
+Claim: “The policy has been a complete disaster for all families.”
+Before: Yes   After: No
+This change follows the newly added rule that exaggeration or unverifiable claims should be judged as False, reducing false positives driven by rhetorical or hyperbolic language (e.g., “complete disaster”).
+Appendix I Case Study: System Prompt Optimization
 To further illustrate the applicability of MAPGD, we present a case study where our method is applied to optimize the system prompt of a large language model assistant.
@@
 Structured Classification and Risk Prioritization Framework
-  1. (1)
+  1. 1.
     Factual Reporting and Descriptive Analysis: Present verified information such as corporate history, operational metrics, and financial data neutrally, before transitioning to evaluative content.
-  2. (2)
+  2. 2.
     Business Analysis: Evaluate performance, market positioning, and strategic initiatives; assess risks by severity, likelihood, and propose contextualized mitigation.
-  3. (3)
+  3. 3.
     Legal Risk Analysis: Examine compliance, regulatory, and contractual risks; assess impact and propose mitigation actions aligned with legal context.
-  4. (4)
+  4. 4.
     Cross-Domain Analysis: For overlapping elements, classify by primary context and document dual-category cases with rationale.
@@
 These findings further support that MAPGD enhances the robustness and effectiveness of system prompts in real-world generation tasks.
-Appendix E Agent Specialization for Different Tasks
+Appendix J Agent Specialization for Different Tasks
 To adapt MAPGD to different domains, we define specialized agent roles tailored to the unique requirements of each task.
 Below are the agent configurations used for the classification, mathematical reasoning, and financial analysis tasks in our experiments.
-E.1. Agent Roles for Classification Tasks
+J.1 Agent Roles for Classification Tasks
 For general classification tasks, the agents focus on the core components of a good prompt: instructions, examples, format, and style.
-instruction_specialist:
+instruction_specialist
 Analyzes and improves task instructions, ensuring clarity, completeness, and executability.
-example_curator:
+example_curator
 Focuses on selecting representative, diverse examples and ensuring consistent formatting.
-format_designer:
+format_designer
 Designs clear output templates and structured formats for better model understanding.
-style_optimizer:
+style_optimizer
 Optimizes language expression for professionalism and task-specific adaptation.
-E.2. Agent Roles for Mathematical Reasoning Tasks
+J.2 Agent Roles for Mathematical Reasoning Tasks
 For mathematical reasoning, the agent roles are adapted to focus on logical decomposition, calculation, and problem interpretation.
-reasoning_specialist:
+reasoning_specialist
 Specializes in enhancing mathematical reasoning processes, ensuring clear step-by-step logic and proper problem decomposition.
-calculation_optimizer:
+calculation_optimizer
 Focuses on improving calculation accuracy, suggesting better computational methods, and ensuring numerical correctness.
-problem_interpreter:
+problem_interpreter
 Specializes in interpreting math word problems, extracting relevant information, and identifying mathematical relationships.
-solution_formatter:
+solution_formatter
 Optimizes mathematical solution presentation, ensures clear formatting and proper answer notation (e.g., #### [final_answer]).
-E.3. Agent Roles for Financial Report Analysis
+J.3 Agent Roles for Financial Report Analysis
 For the complex task of business and financial analysis, agents are given highly specialized roles focusing on data verification, structured report generation, and professional tone.
-instruction_specialist:
+instruction_specialist
 Crafts clear, executable instructions for generating truthful, structured reports and rejecting false or exaggerated inputs.
 Instructions must include steps to verify input data authenticity (e.g., checking financial metrics for realism).
-example_curator:
+example_curator
 Selects and formats diverse examples that demonstrate truthful report generation and false data rejection, including both positive (realistic) and negative (exaggerated) samples.
-format_designer:
+format_designer
 Specializes in designing clear, structured output templates for reports (e.g., using <Business Analysis>, <Financial Report> tags) and error messages (e.g., <Error>).
-style_optimizer:
+style_optimizer
 Optimizes language style and tone for generating professional reports in business, financial, and legal contexts, while ensuring error messages are direct and unambiguous.
-Appendix F Example Prompts Before and After Optimization
+Appendix K Example Prompts Before and After Optimization
 This appendix provides the initial prompts and the final optimized prompts generated by MAPGD for each of the benchmark datasets used in our experiments.
-F.1. LIAR Dataset
-Initial Prompt
-# Task
-Determine whether the Statement is a lie (Yes) or not (No)
-based on the Context and other information.
-# Output format
-Answer Yes or No as labels
-# Prediction
-Text: {text}
-label:
-Optimized Prompt
-# Task
-Your task is to evaluate the given Statement in the context
-of the provided information. If the Statement is found to be
-intentionally false or misleading, it should be classified
-as a lie. In this case, your response should be ’Yes’.
-If the Statement is true or unintentionally false due to a
-mistake or lack of knowledge, it should not be classified
-as a lie. In this case, your response should be ’No’.
-# Output format
-Provide your answer as ’Yes’ or ’No’ labels. ’Yes’ indicates
-that the statement is a lie, while ’No’ indicates that the
-statement is not a lie.
-# Prediction
-Text: {text}
-label:
-F.2. Ethos Dataset
-Initial Prompt
-# Task
-Is the following text hate speech?
-# Output format
-Answer Yes or No as labels
-# Prediction
-Text: {text}
-Label:
-Optimized Prompt
-# Task
-Your task is to analyze the following text and determine if
-it contains hate speech. For the purpose of this task, hate
-speech is defined as any speech, gesture, conduct, writing,
-or display that may incite violence or prejudicial actions
-against or by a protected individual or group, or because it
-disparages or intimidates a protected individual or group.
-Please base your judgment on the explicit content of the text,
-rather than personal interpretations or assumptions.
-# Output format
-Provide your answer in a clear format as either
-’Contains Hate Speech’ or ’Does Not Contain Hate Speech’
-# Prediction
-Text: {text}
-Label:
-F.3. Jailbreak Dataset
-Initial Prompt
-# Task
-Detect if the message is a jailbreak attack,
-i.e. an attempt by a user to break through an
-AI system’s protections
-# Output format
-Answer Yes or No as labels
-# Prediction
-Text: {text}
-Label:
-Optimized Prompt
-# Task
-Classify the message as one of the following labels based on
-its potential to be a jailbreak attack, i.e., an attempt to
-bypass an AI system’s protections:
-- "OvertAttack": The message contains clear, direct attempts
-  to break through system safeguards.
-- "SuspiciousContext": The message shows indirect or
-  contextual cues that may indicate a multi-stage or
-  sophisticated attack attempt.
-- "Benign": The message poses no apparent risk and does not
-  attempt to circumvent protections.
-# Output format
-Select and output exactly one label from the options above.
-# Prediction
-Text: {text}
-Label:
-F.4. GSM8K Dataset
-Initial Prompt
-# Task
-Solve the math word problem step by step.
-# Instructions
-1. Read the problem carefully
-2. Identify what needs to be calculated
-3. Show your work step by step
-4. Provide the final numerical answer
-# Output Format
-Show your reasoning process and end with: #### [final_answer]
-# Problem
-{text}
-Optimized Prompt
-# Task
-Solve the math word problem step by step.
-# Instructions
-1. Read the problem carefully
-2. Identify what needs to be calculated
-3. If there are ambiguous situations in the problem, make
-   reasonable assumptions and state them clearly
-4. Show your work step by step
-5. Provide the final numerical answer
-# Output Format
-Show your reasoning process, including any assumptions made,
-and end with: #### [final_answer]
-# Problem
-{text}
-F.5. AQUARAT Dataset
-Initial Prompt
-# Task
-Solve the math word problem and choose the correct answer
-from the given options.
-# Instructions
-1. Read the problem carefully
-2. Analyze each option
-3. Show your reasoning step by step
-4. Select the correct answer (A, B, C, D, or E)
-# Output Format
-Show your reasoning and end with: Answer: [LETTER]
-(For example: "Answer: A" or "Answer: B")
-# Problem
-{text}
-# Options
-{options}
-Optimized Prompt
-# Task
-Solve the provided problem by systematically classifying its
-type, analyzing all constraints, applying formal mathematical
-reasoning, and verifying the solution against all conditions
-before selecting the correct answer.
-# Instructions
-Adhere strictly to the following structured framework:
-1. **Problem Classification:**
-   * Explicitly identify the core problem domain (e.g.,
-     "Combinatorial Selection," "Constraint Satisfaction,"
-     "Partnership Profit-Sharing," "Work Rate").
-   * Justify the classification with a brief rationale
-     based on the problem statement.
-2. **Constraint Analysis:**
-   * Enumerate all explicit constraints from the problem text.
-   * Infer and list any implicit constraints or logical
-     dependencies.
-   * Categorize each constraint (e.g., Hard/Must-Fulfill,
-     Soft/Optimization) and specify their logical
-     relationships (e.g., AND, OR).
-3. **Mathematical Formalization & Numerical Context
-     Resolution:**
-   * Identify and resolve any ambiguous numerical references
-     (e.g., "monthly" vs. "annual", "together" vs.
+K.1 LIAR Dataset
+Initial Prompt # Task Determine whether the Statement is a lie (Yes) or not (No) based on the Context and other information. # Output format Answer Yes or No as labels # Prediction Text: {text} label:
+Optimized Prompt # Task Evaluate the given Statement using the provided context.
+Classify as Yes if the Statement is intentionally false or misleading (i.e., a lie).
+Classify as No if the Statement is true, or if it is incorrect due to a mistake or lack of knowledge rather than intentional deception.
+# Output format Output exactly one label: Yes or No’.
+Yes indicates the Statement is a lie; No indicates it is not a lie.
+# Prediction Text: {text} label:
+K.2 Ethos Dataset
+Initial Prompt # Task Is the following text hate speech? # Output format Answer Yes or No as labels # Prediction Text: {text} Label:
+Optimized Prompt # Task Your task is to analyze the following text and determine if it contains hate speech.
+For the purpose of this task, hate speech is defined as any speech, gesture, conduct, writing, or display that may incite violence or prejudicial actions against or by a protected individual or group, or because it disparages or intimidates a protected individual or group.
+Please base your judgment on the explicit content of the text, rather than personal interpretations or assumptions.
+# Output format Provide your answer in a clear format as either Contains Hate Speech or Does Not Contain Hate Speech # Prediction Text: {text} Label:
+K.3 Jailbreak Dataset
+Initial Prompt # Task Detect if the message is a jailbreak attack, i.e.
+an attempt by a user to break through an AI systems protections # Output format Answer Yes or No as labels # Prediction Text: {text} Label:
+Optimized Prompt # Task Classify the message as one of the following labels based on its potential to be a jailbreak attack, i.e., an attempt to bypass an AI systems protections: - "OvertAttack": The message contains clear, direct attempts to break through system safeguards.
+- "SuspiciousContext": The message shows indirect or contextual cues that may indicate a multi-stage or sophisticated attack attempt.
+- "Benign": The message poses no apparent risk and does not attempt to circumvent protections.
+# Output format Select and output exactly one label from the options above.
+# Prediction Text: {text} Label:
+K.4 GSM8K Dataset
+Initial Prompt # Task Solve the math word problem step by step.
+# Instructions 1.
+Read the problem carefully 2.
+Identify what needs to be calculated 3.
+Show your work step by step 4.
+Provide the final numerical answer # Output Format Show your reasoning process and end with: #### [final_answer] # Problem {text}
+Optimized Prompt # Task Solve the math word problem step by step.
+# Instructions 1.
+Read the problem carefully 2.
+Identify what needs to be calculated 3.
+If there are ambiguous situations in the problem, make reasonable assumptions and state them clearly 4.
+Show your work step by step 5.
+Provide the final numerical answer # Output Format Show your reasoning process, including any assumptions made, and end with: #### [final_answer] # Problem {text}
+K.5 AQUARAT Dataset
+Initial Prompt # Task Solve the math word problem and choose the correct answer from the given options.
+# Instructions 1.
+Read the problem carefully 2.
+Analyze each option 3.
+Show your reasoning step by step 4.
+Select the correct answer (A, B, C, D, or E) # Output Format Show your reasoning and end with: Answer: [LETTER] (For example: "Answer: A" or "Answer: B") # Problem {text} # Options {options}
+Optimized Prompt # Task Solve the provided problem by systematically classifying its type, analyzing all constraints, applying formal mathematical reasoning, and verifying the solution against all conditions before selecting the correct answer.
+# Instructions Adhere strictly to the following structured framework: 1.
+**Problem Classification:** * Explicitly identify the core problem domain (e.g., "Combinatorial Selection," "Constraint Satisfaction," "Partnership Profit-Sharing," "Work Rate").
+* Justify the classification with a brief rationale based on the problem statement.
+2.
+**Constraint Analysis:** * Enumerate all explicit constraints from the problem text.
+* Infer and list any implicit constraints or logical dependencies.
+* Categorize each constraint (e.g., Hard/Must-Fulfill, Soft/Optimization) and specify their logical relationships (e.g., AND, OR).
+3.
+**Mathematical Formalization & Numerical Context Resolution:** * Identify and resolve any ambiguous numerical references (e.g., "monthly" vs.
+"annual", "together" vs.
 "individual").
-   * Infer and incorporate any implicit numerical variables
-     (e.g., initial quantities, unstated rates, starting
-     points).
+* Infer and incorporate any implicit numerical variables (e.g., initial quantities, unstated rates, starting points).
 * Normalize all units to a consistent basis for calculation.
-   * Define all variables, sets, and parameters using clear
-     notation (e.g., C_total, P_A, S_eligible, nCr).
-   * Structure the problem using appropriate formalisms:
-     timelines, sets, equations, or logical statements.
-   * For combinatorial problems, explicitly state
-     the formula used.
-4. **Solution Execution:**
-   * Perform all calculations step-by-step, showing
-     substitutions and operations.
-   * For combinatorial scenarios, enumerate valid
-     combinations or calculate cardinalities.
-   * Derive the target value (e.g., profit share, number of
-     committees, optimal value).
-5. **Verification & Solution Finalization:**
-   * Systematically verify that the proposed solution
-     satisfies every constraint from Step 2.
-   * For combinatorial answers, confirm counts against
-     constraints using direct checks or complementary
-     counting.
-   * Match the final, verified result to the provided
-     options to select the correct answer.
-# Output Format
-Your final output must follow this structure precisely:
-**Problem Classification:**
-[Your classification and rationale]
-**Constraint
-F.6. SVAMP Dataset
-Initial Prompt
-# Task
-Solve the math word problem step by step.
-# Instructions
-1. Read the problem carefully
-2. Identify what needs to be calculated
-3. Show your work step by step
-4. Provide the final numerical answer
-# Output Format
-Show your reasoning process and end with: #### [final_answer]
-# Problem
-{text}
-Optimized Prompt
-F.7. SVAMP Dataset
-Optimized Prompt
-# Task
-Solve the math word problem through structured, hierarchical
-reasoning. Your primary goal is to correctly parse and compute
-multi-step problems involving distinct object categories and
-their quantitative relationships.
-# Instructions
-1. **Parse and Categorize:**
-   Read the problem and question carefully. Identify and list
-   all distinct object types or categories (e.g., ’packages
-   of gum’, ’boxes of candy’). For each category, extract:
-   - The number of units (e.g., 5 packages).
-   - The quantity per unit, if specified (e.g., 8 pieces
-     per package).
-2. **Structure the Solution Hierarchically:**
-   - **Step 1: Calculate Intra-Category Totals.**
-     For each identified category, calculate its total quantity.
-     If a unit rate is given, perform the multiplication
-     (e.g., ‘5 packages * 8 pieces/package = 40 pieces of gum‘).
+* Define all variables, sets, and parameters using clear notation (e.g., C_total, P_A, S_eligible, nCr).
+* Structure the problem using appropriate formalisms: timelines, sets, equations, or logical statements.
+* For combinatorial problems, explicitly state the formula used.
+4.
+**Solution Execution:** * Perform all calculations step-by-step, showing substitutions and operations.
+* For combinatorial scenarios, enumerate valid combinations or calculate cardinalities.
+* Derive the target value (e.g., profit share, number of committees, optimal value).
+5.
+**Verification & Solution Finalization:** * Systematically verify that the proposed solution satisfies every constraint from Step 2.
+* For combinatorial answers, confirm counts against constraints using direct checks or complementary counting.
+* Match the final, verified result to the provided options to select the correct answer.
+# Output Format Your final output must follow this structure precisely: **Problem Classification:** [Your classification and rationale]
+K.6 SVAMP Dataset
+Initial Prompt # Task Solve the math word problem step by step.
+# Instructions 1.
+Read the problem carefully 2.
+Identify what needs to be calculated 3.
+Show your work step by step 4.
+Provide the final numerical answer # Output Format Show your reasoning process and end with: #### [final_answer] # Problem {text}
+Optimized Prompt # Task Solve the math word problem through structured, hierarchical reasoning.
+Your primary goal is to correctly parse and compute multi-step problems involving distinct object categories and their quantitative relationships.
+# Instructions 1.
+**Parse and Categorize:** Read the problem and question carefully.
+Identify and list all distinct object types or categories (e.g., packages of gum’, boxes of candy’).
+For each category, extract: - The number of units (e.g., 5 packages).
+- The quantity per unit, if specified (e.g., 8 pieces per package).
+2.
+**Structure the Solution Hierarchically:** - **Step 1: Calculate Intra-Category Totals.** For each identified category, calculate its total quantity.
+If a unit rate is given, perform the multiplication (e.g., ‘5 packages * 8 pieces/package = 40 pieces of gum‘).
 If no unit rate exists, use the given quantity directly.
-   - **Step 2: Perform Inter-Category Operations.**
-     Using the totals from Step 1, now perform the final operation
-     as required by the question (e.g., sum all category totals
-     for a grand total, or find the difference between two
-     category totals for a comparison).
-3. **Show Your Work:**
-   Present your reasoning clearly, reflecting this two-step
-   hierarchical process. First, show all calculations within
-   categories. Then, show the final combination of these
-   category totals.
-4. **Final Answer:**
-   Box your final numerical result.
-# Output Format
-Reasoning: [Your step-by-step reasoning here]
-#### [final_answer]
-# Problem
-{text}
-# Question
-{question}
+- **Step 2: Perform Inter-Category Operations.** Using the totals from Step 1, now perform the final operation as required by the question (e.g., sum all category totals for a grand total, or find the difference between two category totals for a comparison).
+3.
+**Show Your Work:** Present your reasoning clearly, reflecting this two-step hierarchical process.
+First, show all calculations within categories.
+Then, show the final combination of these category totals.
+4.
+**Final Answer:** Box your final numerical result.
+# Output Format Reasoning: [Your step-by-step reasoning here] #### [final_answer] # Problem {text} # Question {question}